### v1

# GHP-MOFassemble: Diffusion modeling, high throughput screening, and molecular dynamics for rational discovery of novel metal-organic frameworks for carbon capture at scale

Abstract
We introduce GHP-MOFassemble, a Generative artificial intelligence (AI), High Performance framework to accelerate the rational design of metal-organic frameworks (MOFs) with high CO2 capacity and synthesizable linkers. Our framework combines a diffusion model, a class of generative AI, to generate novel linkers that are assembled with one of three pre-selected nodes into MOFs in a primitive cubic (pcu) topology. The CO2 capacities of these AI-generated MOFs are predicted using a modified version of the crystal graph convolutional neural network model. We then use the LAMMPS code to perform molecular dynamics simulations to relax the AI-generated MOF structures, and identify those that converge to stable structures, and maintain their porous properties throughout the simulations. Among 120,000 pcu MOF candidates generated by the GHP-MOFassemble framework, with three distinct metal nodes (Cu paddlewheel, Zn paddlewheel, Zn tetramer), a total of 102 structures completed molecular dynamics simulations at 1 bar with predicted CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of hMOFs in the hypothetical MOF (hMOF) dataset in the MOFX-DB database. Among these candidates, 18 have change in density lower than 1% during molecular dynamics simulations, indicating their stability. We also found that the top five GHP-MOFassemble’s MOF structures have CO2 capacities higher than 96.9% of hMOF structures. This new approach combines generative AI, graph modeling, large-scale molecular dynamics simulations, and extreme scale computing to open up new pathways for the accelerated discovery of novel MOF structures at scale.
  • July 13, 2023
Keywords: diffusion model, pre-trained regression model, metal-organic framework, carbon capture
1 Introduction
Metal-organic frameworks (MOFs) have garnered much research interest in recent years due to their diverse industrial applications, including gas adsorption and storage [1], catalysis [2], and drug delivery [3]. As nanocrystalline porous materials, MOFs are modular in nature [4], with a specific MOF defined by its three types of constituent building blocks (inorganic nodes, organic nodes, and organic linkers) and topology (the relative positions and orientations of building blocks). MOFs with different properties can be produced by varying these building blocks and their spatial arrangements. Nodes and linkers differ in terms of their numbers of connection points: inorganic nodes are typically metal oxide complexes whereas organic nodes are molecules, both with three or more connection points; organic linkers are molecules with only two connection points.
MOFs have been shown to exhibit superior chemical and physical CO2 adsorption properties. They can be recycled in appropriate operating environments for varying numbers of times before undergoing significant structural degradation. However, their industrial applications have not yet reached full potential due to stability issues, such as poor long-term recyclability [5], and high moisture sensitivity [6]. For instance, the presence of moisture in adsorption gas impairs a MOF’s CO2 capture performance, which may be attributed to MOFs having stronger affinity toward water molecules than to CO2 molecules [7]. Other investigations of MOF stability issues [8, 9, 10] have shown [11, 12] that the gas adsorption properties of MOFs can be enhanced by tuning the building blocks used. Yet progress has been difficult due to the enormous chemical space of building blocks that makes exhaustive search with traditional experimental methods impractical [13].
Here we propose GHP-MOFassemble, a novel high-throughput computational framework to accelerate the discovery of MOF structures with high CO2 capacities and synthesizable linkers. To demonstrate the utility of our framework, we apply it to the rational design of MOFs with pcu topology and three types of inorganic nodes: Cu paddlewheel, Zn paddlewheel, and Zn tetramer. The key component of our generative framework is a diffusion model, Difflinker [14], which enables efficient generation of new linker molecules from a set of molecular fragments.
Previous attempts to discover useful MOFs have proceeded via high-throughput screening of existing datasets [15, 16, 17], an approach that necessarily limits the search to known MOFs. Our GHP-MOFassemble framework instead probes the MOF design space by employing a molecular generative diffusion model (DiffLinker [14]) to generate chemically diverse and unique MOF linkers, which it assembles with pre-selected metal nodes to form novel MOF structures. It then screens those structures with a pre-trained regression model, a modified version of CGCNN [18], to identify high-performing MOF candidates for carbon capture.
2 Related Work
Previous efforts in the search for new MOF structures with exceptional gas adsorption properties include high-throughput screening and generative modeling. High-throughput screening approaches comprise both database search and machine learning (ML)–assisted screening.
Database search methods. These methods apply filters to identify attractive candidates in large databases of MOF structures plus calculated properties obtained by using molecular simulations. For example, Li et al. [16] search the computation-ready, experimental MOF database (CoRE DB) for MOFs with high $\textrm{CO}_{2}$ uptake when moisture is present, and Altintas et al. [19] search both CoRE DB and the Cambridge Structural Database non-disordered MOF subset for MOFs with high $\textrm{CH}_{4}/\textrm{H}_{2}$ selectivity.
ML–assisted screening. Database search methods require expensive molecular simulation calculations for every target molecule. ML-assisted screening avoids this difficulty by using a regression model trained on a smaller training dataset to predict target properties for a large number of new test structures. This approach has been widely applied to gas adsorption and separation property predictions of MOFs, including CO2/H2 separation [20] and CH4 adsorption [21]. ML-assisted methods commonly make use of both geometrical features (e.g., dominant pore size, void fraction [22]) and chemical features (e.g., atom type, electronegativity [23]). A compound feature, the atomic property weighted radial distribution function (AP-RDF) [24] has been shown to improve regression model performance in finding MOF structures with high CO2 capacity [25]. This feature is constructed by using both chemistry and local geometry of atomic sites in MOF structures. As an alternative to the use of hand-engineered machine learning features, neural networks have also been applied to MOF research. In a recent study, the Atomistic Line Graph Neural Network (ALIGNN), trained on the hMOF dataset [26], was applied to search CoRE DB for high-performing MOF candidates for carbon capture [27].
Generative modeling. In the generative modeling approach, candidate molecules are not drawn from a database but are generated de novo via methods that produce novel compounds that are expected to have a desired set of chemical features. One notable example is the recently proposed Supramolecular Variational Autoencoder (SmVAE[28], which uses a semantically constrained graph-based canonical code (RFcode) to encode MOF building blocks. The variational autoencoder framework structure allows this model to interpolate between existing MOF structures and enables isoreticular optimization of MOF structures toward higher CO2 capacity and CO2/N2 selectivity.
This work. Proposed generative models include not only variational autoencoders (VAEs) but also generative adversarial networks (GANs), normalizing flows, autoregressive models, and diffusion models [29]. In this work, we adopt a diffusion model named DiffLinker to generate novel MOF linkers. Diffusion models use a probability distribution and Markovian properties to generate new data via a denoising step. First, Gaussian noise is added to the input samples to yield noisy data. A neural network is then trained to learn what noise was added. The trained network is then used to reversely transform (i.e., denoise) the noisy data back into target samples that resemble molecules from the training data distribution. This method has been widely applied to drug discovery to speed up the design of new ligands and ligand-protein docking configurations [30, 31, 32, 33].
Two broad categories of molecular representation schemes have been used in diffusion model-based ligand generators: molecular graphs and 3D coordinates. In the former case (Digress/Congress [34] is an example), atoms are represented as nodes and bonds as edges. In the latter case, the 3D atomic coordinates of molecules are generated directly, as in DiffLinker [14] and E(3) equivariant diffusion model (EDM[35]. Given the success of diffusion models in drug discovery [32, 30, 33, 36, 37], in Section 3.1 we demonstrate how to transfer the idea to the iso-reticular design of MOF structures by varying MOF linkers while fixing node and topology. The diffusion model is used specifically for MOF linker design, as we describe in Section 3.2.
3 Methods
Our GHP-MOFassemble framework has three components:
  • Decompose. We use a molecular fragmentation algorithm to decompose the MOF linkers found in high-performing MOFs within the hMOF dataset [26]—an open source dataset that provides, for each of 137,652 hypothetical MOF structures, its MOFid, MOFkey, geometric features, and isotherm data for six adsorption gases (CO2, N2, CH4, H2, Kr, Xe) at 0.01, 0.05, 0.1, 0.5, and 2.5 bar—covering the pressure ranges of cyclic adsorption gas separation processes in industrial applications. The adsorption properties of gases provided in this dataset were calculated using Grand Canonical Monte Carlo (GCMC) calculations [26], as described in Wilmer et al. [38].
  • Generate. We use a diffusion model to generate new MOF linkers. We then screen the generated linkers by removing linkers with S, Br and I elements (we call this step “element filter” in later text), and evaluate their quality using five different scores to quantify their synthesizability, validity, uniqueness, and internal diversity. Linkers that passed the element filter are then assembled with one of three pre-selected nodes into new MOF structures in the pcu topology.
  • Predict. We use a pre-trained regression model to predict the CO2 capacities of the newly generated MOF structures.
In the following sections we describe each of these components in detail.
3.1 Decomposing MOF linkers into molecular fragments
The first component of the GHP-MOFassemble framework, Decompose, decomposes linkers from high-performing MOF structures in the hMOF dataset into their molecular fragments. As illustrated in Figure 1, this process applies the following three steps to a specified node-topology pair, which in the figure is the most frequent node-topology pair in the hMOF dataset, Zn tetramer-pcu. Note that for the pcu topology, one linker is orientated along each of the x, y, and z directions, and thus at most three linker types are possible.
  1. 1.
    Select: We select high-performing MOFs with the given node-topology pair from hMOF. We define here a high-performing MOF structure as one with CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of CO2 capacity.
  2. 2.
    Extract: We extract the linkers Simplified Molecular-Input Line-Entry System (SMILES) strings for each high-performing MOF identified in step (i), eliminate those for MOFs with more than three unique linker types, and assemble those that remain into a linker dataset.
  3. 3.
    Fragment: We fragment the linkers produced in step (ii) to obtain their chemically relevant fragment-connection atom pairs, which we assemble into a molecular fragment dataset.
The extraction of linkers in step (ii) is straightforward because the MOFid provided by hMOF for each MOF specifies the SMILES strings of its constituent inorganic and organic building blocks, including its linkers, as well as a format signature and a topology code [39]. Together, these elements uniquely define the topology and building blocks of a given MOF structure.
Step (iii) use the Matched Molecular Pairs Algorithm (MMPA) [40] as implemented in DeLinker [41] to generate molecular fragments of a given molecule by breaking chemical bonds between atom pairs. We set the minimum number of connection atoms, the minimum fragment size, and the minimum path length to 3, 5, and 2, respectively. Moreover, we only consider the case where the fragments are at least two atoms away from each other. The chemically relevant fragment-connection atoms pairs are then used to form the molecular fragment dataset.
Figure 1: Schematic representation of Decompose, the first component of our GHP-MOFassemble framework. It consists of three steps, which we showcase for Zn tetramer-pcu MOFs. First, the Select step selects the Zn tetramer-pcu MOFs in hMOF that are high-performing. Then the Extract step extracts unique linkers from those MOFs. Finally, the Fragment step generates unique linker fragments. The color scheme of elements is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white.
3.2 Generating new MOF structures
The Generate component employs the pre-trained diffusion model DiffLinker [14] to generate new MOF linkers, and then assembles those new linkers with one of three pre-selected nodes into new MOF structures in the pcu topology. It comprises three steps: Diffuse and Denoise; Screen and Evaluate; and Assemble. An example of these three steps for Zn tetramer-pcu MOFs is shown in Figure 2.
3.2.1 Diffuse and Denoise.
We apply the pre-trained diffusion model DiffLinker [14] to the molecular fragments output by the Decompose component to generate novel MOF linkers. This model connects the molecular fragments supplied as input (also known as context) with a specified number of sampled atoms. We vary the number of sampled atoms from five to 10, a range chosen because we found that with fewer than five sampled atoms, straight chains or branched chains may be obtained, while with more than 10, ring structures may be present. We found that the range of five to 10 sampled atoms resulted in novel and chemically diverse MOF linkers.
DiffLinker applies a denoising process to determine the atomic species and Cartesian coordinates of sampled atoms by using a decoder neural network architecture named E(3)-Equivariant Graph Neural Network (EGNN[42]. This graph neural network predicts zero-mean centered Cartesian coordinates noise and one-hot encoding noise of atomic species, and accounts for equivariance due to translation, rotation, and reflection of molecules, i.e., Euclidean group 3 (E(3)). Since the pre-trained DiffLinker model was trained on the GEOM dataset [43], it enables sampling of connection atoms with high chemical and structural diversity.
DiffLinker outputs the 3D coordinates and the atomic species of heavy atoms in the linker molecules, rather than SMILES strings or 2D graphs. Since hydrogen atoms are implicit in the DiffLinker model, their 3D coordinates are not output by the model. To generate the spatial configurations of all-atom molecules, we employ openbabel to convert the DiffLinker generated molecules to SMILES strings, which in turn are used to generate the 3D configurations of linkers with explicit hydrogen atoms. Openbabel uses distance-based heuristics to determine bond connectivity in a given molecule [44], which is critical to the assignment of the number of hydrogen atoms. After the hydrogen addition step, we identify the dummy atoms which contains information about how the linkers connect with metal nodes.
The diffusion process involves two consecutive steps. The first adds Gaussian noise to the original data (i.e., x), yielding noisy input (i.e., $z_{t}$ at time t). Then, the denoising step applies the neural network-based noise removal operation. Mathematically, the following Markovian properties and equations are satisfied:
$q(z_{t}|z_{t-1})$ $=$ $\mathcal{N}(z_{t};\bar{\alpha_{t}}z_{t-1},\bar{\sigma}_{t}^{2}\textbf{{I}})\,,$ (1)
$q(z_{t}|\textbf{x})$ $=$ $\mathcal{N}(z_{t};\alpha_{t}\textbf{x},\sigma_{t}^{2}\textbf{{I}})\,,$ (2)
$p(z_{t-1}|z_{t})$ $=$ $q(z_{t-1}|\textbf{x},z_{t})\,,$ (3)
$q(z_{t-1}|\textbf{x},z_{t})$ $=$ $\mathcal{N}(z_{t-1};\mu_{t}^{\theta}(\textbf{x},z_{t};\alpha_{t},\sigma_{t}),\xi_{t}\textbf{{I}})\,,$ (4)
$\mu_{t}^{\theta}(\textbf{x},z_{t};\alpha_{t},\sigma_{t})$ $=$ $A_{t}z_{t}+B_{t}\textbf{x}\,,$ (5)
$\hat{\textbf{x}}=\textbf{x}=C_{t}z_{t}$displaystyle-- $D_{t}\epsilon^{\theta}_{t}(z_{t},t)\,,$ (6)
[…]
Equations (1) and (2) represent a diffusion process that is conditioned on the previous value, $z_{t-1}$, and initial value, x, to predict a current value, $z_{t}$. The denoising processes are described by Equations (3) and (4), which predict a previous value conditioned on (either real or predicted) initial value and current value. Equations (5) and (6) provide a detailed evolution for $\mu^{\theta}_{t}$ which represents the learned denoising mean parameterized by $\theta$, as well as x, a real initial value, approximated by $\hat{\textbf{x}}$. The training objective is to learn a neural network, i.e., EGNN, to predict a denoising value so that random noise can be denoised to a physically and chemically valid molecule during the generative process. In practice, we predict $\epsilon^{\theta}_{t}$ (learned noise value) rather than $\hat{\textbf{x}}$ directly (i.e., Equations (5) and (6)) for better prediction [45]. Since DiffLinker generates chemically valid molecules, EGNN needs to take more regularization into account, such as E(3) and O(3) (i.e., orthogonal group in dimension 3: rotations and reflections). Thus, invariant features such as one-hot and equivariant features such as atomic coordinates updates need to be considered [14].
3.2.2 Screen and Evaluate.
To ensure consistency of elements in hMOF linkers and generated linkers, we manuallyfilter out linkers that contain elements not present in the hMOF dataset. Since the pre-trained DiffLinker model we used in this work was trained on the GEOM dataset, a total of nine heavy elements may be sampled, including C, N, O, F, P, S, Cl, Br, and I. Among these elements, S, Br, and I do not appear in the hMOF dataset, therefore we remove linkers that contain these three elements. For each molecular fragment, we then perform sampling 20 times. Each sampling step may yield a different set of molecules because of the random nature of the denoising process. The probabilistic nature of the diffusion model allows it to generate linkers from an extensive linker design space beyond that of the hMOF dataset.
We then use five metrics to evaluate the quality of the remaining linkers. The first two metrics are commonly used heuristic measures of synthesizability: the synthetic accessibility score (SAscore), and the synthetic complexity score (SCscore) [46].
The SAscore, as defined by Ertl and Schuffenhauer [47], is based on analysis of one million PubChem molecules, and combines fragment contributions from molecule substructures with a complexity penalty that accounts for molecular size and for structural features of molecules such as the presence of rings. The SCscore is computed by using a neural network trained on 12 million chemical reactions from the Reaxys database to estimate the number of reaction steps required to produce a molecule [46]. For both SAscore and SCscore, the higher the values are, the more difficult it is to synthesize the linker, hence less desirable.
To evaluate the capability of GHP-MOFassemble to generate a valid, novel, and diverse set of linkers, we leverage the MOSES framework [48] to compute three additional metrics: the fraction with valid SMILES strings (validity), the fraction that are unique (uniqueness), and their dissimilarity (internal diversity). We compute the fraction of linkers that are unique by comparing their canonical SMILES strings, and use their average pairwise molecular Tanimoto dissimilarity [49] as measures of diversity. We compute the diversity of the generated linker set by using two internal diversity scores $(\textrm{IntDiv}_{1}(G),\textrm{IntDiv}_{2}(G))$. We present the results of these metrics in Section 4.2.
Figure 2: The Generate component of the GHP-MOFassemble framework involves Diffuse and Denoise, Screen, and Assemble steps, as shown here for Zn tetramer-pcu MOFs. First, we generate new linkers via the Diffuse and Denoise step. We then add hydrogen atoms to ensure correct valency of the generated linkers. Next, we identify dummy atoms by replacing the two carbon atoms in the carboxyl groups of each generated linker with dummy atoms. Linkers with dummy atoms then undergo the Screen step, which removes linkers with S, Br and I. These three elements are present in the GEOM dataset but not in hMOF dataset, and thus we filter out linkers with these three elements. As a result, this step reduces the number of potential linkers to 5,162. The generated linkers’ molecular statistics are quantified using five metrics, including SCscore, SAscore, validity, uniqueness, and internal diversity. Finally, in Assemble step, we build 40,000 new MOFs with the Zn tetramer node in pcu topology. As before, the color code used for the atoms shown is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white.
3.2.3 Assemble.
Once we have generated linkers that meet the five metrics mentioned above, we can assemble them by connecting the node and linker building blocks. To construct MOFs, we need to guide this assembly process. We do that by using dummy atoms, which indicate the points where the building blocks are to be connected. In practice, this is done as follows. Our parsing of MOFids generates, for MOFs with Zn tetramer nodes, linkers that carry two carboxyl groups, which by definition are part of metal nodes instead of linkers. Such wrong assignment of carboxyl groups is due to how the MOFid algorithm parses MOF structures. To reflect the correct molecular structure of linkers, the two carbon atoms in the carboxyl groups are identified as dummy atoms, and the redundant four oxygen atoms and two hydrogen atoms are removed. An illustration of dummy atom identification for linkers containing carboxyl groups is shown in the left panel of Figure 11.
For MOFs with Cu paddlewheel and Zn paddlewheel node, however, another type of linker containing heterocyclic rings exists. For this type of linker, the two atoms that are nitrogen-metal bond distance away from the terminal nitrogen atoms on the heterocyclic rings are identified as dummy atoms. After the dummy atoms are correctly identified, three randomly selected linkers (duplicates allowed) are assembled with one of the three pre-selected nodes into MOFs in the pcu topology. An illustration of dummy atom identification of linkers containing heterocyclic groups is shown in the right panel of Figure 11.
More than half of the hMOF structures are catenated MOFs, i.e., MOFs with interpenetrated lattices. By varying the level of interpenetration, it is possible to generate MOFs with different pore sizes, with a higher catenation level generally corresponding to smaller pores. We denote the four catenation levels in hMOF, with increasing number of interpenetrated lattices, as cat0, cat1, cat2 and cat3. To generate MOFs with high structural diversity, we applied the site translation method as implemented in Pymatgen [50] to generate MOFs with different catenation levels. For catenated MOF structures, we keep the spacing between interpenetrated lattices the same, so as to ensure equal pore sizes.
3.3 Applying regression model to estimate MOF $\textrm{CO}_{2}$ capacity
The Predict component of the GHP-MOFassemble framework uses a regression model to estimate the CO2 capacity of newly generated pcu MOF structures. We use for this purpose a modified version of the Crystal Graph Convolutional Neural Network (CGCNN) model, developed in our previous work [18], which uses an adjacency list to format node and edge embeddings, rather than the adjacency matrix format of the original CGCNN, a change that enhances model training speed, training stability, and prediction accuracy. We trained this modified CGCNN model on all hMOF structures, with a target property of CO2 capacity at 0.1 bar.
We create our ensemble regression model by training the modified CGCNN model three times and averaging the results. We applied the resulting model to the assembled pcu MOF structures produced by the Generate step of Section 3.2 to identify those with a predicted CO2 capacity higher than 2 mmol/g. These MOFs (see Section 4.4.2) are candidates for further investigation by, for example, by performing CO2 capacity validation using Grand Canonical Monte Carlo (GCMC) simulations.
4 Results
By combining the generative power of diffusion models and the computing power of supercomputers, the GHP-MOFassemble framework enables rapid generation and assembly of chemically diverse MOF structures. Using the hMOF dataset as a starting point, our framework accelerates the computational design of high-performing MOF structures for carbon capture by high-throughput screening of novel MOF structures with high CO2 capacities. Our key findings are summarized below.
4.1 Analysis of the hMOF dataset
The three most frequent node-topology pairs in the hMOF dataset, accounting for around 74% of its MOFs, are the Cu paddlewheel-pcu, Zn paddlewheel-pcu, and Zn tetramer-pcu structures shown in the left panel of Figure 3. As summarized in Table 1, Zn tetramer-pcu is the most abundant, followed by Cu paddlewheel-pcu and then Zn paddlewheel-pcu. The cumulative distribution functions of the CO2 capacities of these three types of MOF structure are shown in the right panel of Figure 3. Only the 78,238 MOFs from hMOF with correctly parsed MOFids and valid SMILES were used for analysis. We gather the molecular fragments numbers in Table 1, produced through the Fragment step described in Section 3.1. In Table 1 high-performing MOFs with three parsed linkers were selected, and their unique linkers parsed by using the MMPA (Matched Molecular Pairs Algorithm) algorithm.
In Table 1, the number of output molecular fragment conformers (last column) is much less than the number of unique linker SMILES (second to last column) for two reasons. First, around 56% of linkers did not pass the valency check, which may be due to the intrinsic limitations in the parsing of MOF linkers.
Second, around 90% of linkers that passed the valency check share similar molecular fragments, and thus many duplicated molecular fragments exist. The successfully parsed molecular fragment conformers are subsequently used for linker generation.
Figure 3: Depictions of the most frequent node-topology pairs in hMOF (left panel) and cumulative distribution functions of the 0.1 bar CO2 capacities (right panel).
Table 1: Data on the hMOF with most frequent node-topology pairs. PW, TM, and HP-MOF stand for paddlewheel, tetramer, and high-performing MOF, respectively.
node-topology total MOFs HP-MOFs HP-MOFs with unique linker unique molecular
three linkers SMILES fragment conformers
Cu PW-pcu 29,714 1,458 1,016 3,330 180
Zn PW-pcu 28,529 1,314 834 3,221 162
Zn TM-pcu 43,874 2,129 1,388 3,265 198
Figure 4: Empirical cumulative distribution functions of the 0.1 bar CO2 capacities of hMOF structures at different catenation levels. The x axis is capped at 4 mmol/g to preserve details.
Figure 4 shows the empirical cumulative distribution functions of the CO2 capacities of hMOF structures at different catenation levels. Therein, we observe that a higher percentage of catenated MOFs (cat1, cat2 and cat3) are high-performing, as compared to the uncatenated MOFs (cat0). This result confirms that catenation is an important factor when designing new MOF structures with high CO2 capacity. In Table 5, we summarize the amount of origin shift for each of the four levels of catenation.
Figure 5 shows the pairwise relationships of the CO2 capacities of hMOF structures at the five pressures. We observe a strong correlation of CO2 capacities between 0.05 bar and 0.1 bar, with a Pearson’s correlation coefficient of 0.86. For other pairs of pressures, the CO2 capacities are only weakly correlated, with a decreasing correlation for larger pressure differences. Moreover, the distributions of CO2 capacities at all pressures exhibit long tails at high value ranges, which indicate that the majority of MOFs are low-performing and high-performing MOFs are uncommon.
Figure 5: Pairplot of the CO2 capacities of hMOF structures at different adsorption pressures. The x- and y-axis are adsorption pressures. Different colors indicate MOFs with different node-topology pairs.
4.2 Linker generation and evaluation
For all three MOF types, we start with 540 molecular fragments extracted from high-performing hMOF structures, and use DiffLinker to generate a large pool of 64,800 linkers with the number of sampled connection atoms ranging from 5 to 10. For each linker, sampling is performed 20 times, as described in Section 3.2. Since these linkers only contain heavy atoms, to generate all-atom linker molecules, we apply openbabel to add hydrogen atoms, which results in 56,257 linkers. Next, dummy atom identification is performed to generate information that enables assembly with metal nodes. A total of 16,162 linkers with generated dummy atoms are generated. These linkers are passed through the element filter, which removes linkers that contain S, Br and I elements, further reducing number to 12,305 linkers. The details of the number of linkers in each step are summarized in Table 7.
We show in Figure 6 the synthetic accessibility score (SAscore) and synthetic complexity score (SCscore) values, defined in Section 3.2.2, for the remaining linkers. We observe that as the number of sampled atoms increases, both distributions generally shift to the right, with the exception of the SAscore distribution with six sampled atoms, which is to the left of that with five atoms. The general trend indicates that linkers become harder to synthesize as the molecules become bigger. This result is expected because as more atoms are sampled, more complex substructures may be present. We note that no linker has zero or very large SAscore or SCscore, values that would indicate unsynthesizability.
Figure 6: Distributions of synthetic accessibility score (SAscore, top row) and synthetic complexity score (SCscore, bottom row) of DiffLinker-generated MOF linkers with dummy atoms. The subfigure titles are node-topology pairs, and the colors listed in the legend indicate the number of sampled atoms, from five to 10 inclusive.
We show in Table 2 the validity, uniqueness, and internal diversity metrics, defined in Section 3.2.2, grouped by number of sampled atoms and node-topology pairs of the corresponding MOFs. The validity column confirms that all generated linkers are valid. The last three columns, when reviewed from top to bottom, reveal a considerable increase in linker uniqueness and a more modest increase in internal diversity as the number of sampled atoms increases. High uniqueness values indicate that our model is generating non-duplicate molecules, whereas high internal diversity values indicate that the generated molecules are chemically diverse. Internal diversity (which indicates how similar/dissimilar a specific linker is to the rest of the population) is computed using two internal diversity scores: IntDiv1 and IntDiv2 [48], also shown in Table 2—see C for details on how these scores are computed. The increase in these metrics as more atoms are sampled is expected because the degrees of freedom of atomic species and their spatial coordinates increase with molecular size.
Table 2: Statistics of the generated linkers that correspond to different MOF types (first column), number of sampled atoms (${\cal{N}}$ in the second column), number of carboxyl linkers (${\cal{C}}$, fourth column), and number of heterocyclic linkers (${\cal{H}}$, fifth column). In total, there are 12,305 linkers with identified dummy atoms.
[…]
unique
$\rm IntDiv_{1}$ $\rm IntDiv_{2}$
Cu paddlewheel-pcu 5 1,240 774 466 1
0.708
0.789 0.769
Zn paddlewheel-pcu 5 1,184 775 409 1
0.692
0.783 0.762
Zn tetramer-pcu 5 1,666 1,666 0 1
0.734
0.739 0.721
Cu paddlewheel-pcu 6 1,117 761 356 1
0.764
0.819 0.796
Zn paddlewheel-pcu 6 992 681 311 1
0.761
0.809 0.787
Zn tetramer-pcu 6 1,532 1,532 0 1
0.766
0.749 0.732
Cu paddlewheel-pcu 7 499 383 116 1
0.958
0.836 0.820
Zn paddlewheel-pcu 7 453 329 124 1
0.929
0.829 0.812
Zn tetramer-pcu 7 709 709 0 1
0.958
0.782 0.770
Cu paddlewheel-pcu 8 330 245 85 1
0.992
0.835 0.822
Zn paddlewheel-pcu 8 338 247 91 1
0.978
0.831 0.817
Zn tetramer-pcu 8 550 550 0 1
0.983
0.791 0.780
Cu paddlewheel-pcu 9 281 191 90 1
0.99
0.837 0.824
Zn paddlewheel-pcu 9 257 175 82 1
0.993
0.834 0.821
Zn tetramer-pcu 9 389 389 0 1
0.986
0.794 0.784
Cu paddlewheel-pcu 10 235 165 70 1
0.996
0.842 0.830
Zn paddlewheel-pcu 10 217 148 69 1
0.996
0.837 0.826
Zn tetramer-pcu 10 316 316 0 1
0.995
0.798 0.789
4.3 MOF assembly
We generate new MOF structures by assembling three randomly selected DiffLinker-generated linkers with one of the three most frequent nodes in the hMOF dataset. Random sampling of linkers ensures that the selected linkers cover a large design space. For each node-linker-topology combination, we considered four levels of catenation (cat0, cat1, cat2, cat3). We generated a total of 120,000 MOFs as follows: we have four catenation levels and three node candidates. The random sampling of 10,000 linkers for each catenation level-node candidate pair generates 120,000 total MOFs of different catenation-node-linker combinations. We provide more details in Section 3.2.3 on how these new MOFs were assembled.
4.4 Regression model for MOF $\textrm{CO}_{2}$ capacity prediction
4.4.1 Pre-training on hMOF structures
We employ the modified version of the CGCNN model [18] as a predictive model of the CO2 capacity of MOF structures. We trained this model on 90% of the MOF structures in hMOF, along with their CO2 capacities at 0.1 bar, leaving 10% as a holdout set for testing. We trained the modified version of the CGCNN model three times to create an ensemble of models in order to mitigate biases in individual models. We summarize in Table 6 and Figure 12 the prediction errors of each model on the 10% test set and the distribution of standard deviations, respectively.
Looking next at the ensemble model, we show in the left panel of Figure 7 its predictions for the 10% test set. The ensemble model has a mean absolute error (MAE) of 0.093 mmol/g on the test set. Using CO2 capacity of 2 mmol/g as a threshold, we repurposed our predictive model as a classifier to categorize MOFs into low and high performers (with predicted CO2 capacities below and above the threshold, respectively) and high performers. Using this scheme, the confusion matrix for identifying low and high performers is shown in the right panel of Figure 7.
We conclude from the confusion matrix that the pre-trained model classifies both low and high performers with high accuracy, with 98.4% (13551 out of 13765) of test samples correctly classified. As the test set is heavily imbalanced, with many more low performers than high performers, we also calculated the balanced accuracy of classification [51] (the sum of true positive rate and true negative rate divided by 2), obtaining a value of 90.7%. The majority of the 214 misclassified samples lie close to the decision boundaries, as shown in the red and purple regions of the left panel of Figure 7. Therefore, we also conclude that the ensemble model is capable of differentiating low and high performers.
Figure 7: Left panel, predictive performance of the ensemble model on the 10% test set. Right panel, confusion matrix of the ensemble model in classifying low and high performers in the test set. The classification threshold is 2 mmol/g.
4.4.2 Predictions for newly generated MOF structures
When using the ensemble model to infer the CO2 capacities of newly generated MOF structures, we employ the average of the predictions made by the three independent models as the predicted capacity. In addition, we filter out structures for which the standard deviation of the predicted values obtained from the three models is larger than 0.2 mmol/g.
We treat the standard deviation of predictions from the three models as a measure of the uncertainty of model predictions, which we view as arising from the model’s difficulty in learning certain data points. We chose the threshold for the standard deviation filter of 0.2 mmol/g because it is sufficiently small (with 96% of data points below the threshold) that for low CO2 capacity predictions, the predicted values across the ensemble are similar. On the other hand, for high CO2 capacity predictions, it also ensures that outlier values (e.g., predictions made by the three models are vastly different) are filtered out, therefore minimizing the overall error.
We show in Figure 8 the empirical cumulative distribution functions of both the hMOF (dashed lines) and predicted (sold lines) CO2 capacities. We observe that there are more high-performing cat2 and cat3 MOFs among the generated structures than in the hMOF dataset, as indicated by lower cumulative density values at 2 mmol/g (black dashed line). On the other hand, there are fewer high-performing cat0 and cat1 MOFs, as indicated by higher cumulative density values at 2 mmol/g. This result indicates that within the generated MOF design space, there are more high-performing candidates with higher catenation levels compared to lower catenation levels.
Figure 8: Comparison of empirical cumulative distribution functions of the predicted CO2 capacities of generated and hMOF MOFs at different catenation levels.
Out of the 120,000 hypothetical MOF structures (see Section 4.3 for MOF assembly details), a total of 6,020 were predicted to be high-performing MOFs with CO2 capacity higher than 2 mmol/g at 0.1 bar. Categorization of the predicted high-performing MOFs by node-topology pair and catenation level (see Table 3) shows that the majority of Cu paddlewheel-pcu and Zn tetramer-pcu are cat2 and cat3, whereas the majority of high-performing Zn paddlewheel-pcu are cat1 and cat2. Table 4 shows the building blocks and catenation levels in the top five pcu MOF candidates. For linkers, the corresponding molecules (without dummy atoms) are shown for ease of visualization. Note that some substructures appear more often than others. For example, ring structures appear in most of the linkers in the top five candidates, and many of the DiffLinker-generated molecular substructures (in between the terminal fragments) also contain ring structures. The frequent occurrence of rings in linkers of high performing MOFs may be due to the presence of aromatic linkers, as reported by phenylene et al. [52].
We compared the chemical similarity of the linkers in predicted high-performing MOF structures with those in high-performing hMOF structures. As shown in Figure 13, we find that most linkers in the predicted high-performing MOF structures are vastly different from those in hMOF, indicating that the GHP-MOFassemble framework generate novel MOF structures with chemically unique linkers.
Table 3: Numbers of predicted high-performing MOFs found in the hMOF dataset.
node-topology cat0 cat1 cat2 cat3
Cu paddlewheel-pcu 0 170 1,184 628
Zn paddlewheel-pcu 9 789 929 395
Zn tetramer-pcu 0 3 794 1,119
Table 4: Details of the five pcu MOF candidates generated by GHP-MOFassemble with the highest predicted CO2 capacities. Their stability was confirmed using molecular dynamics simulations, and their 3D structures are shown in Figure 10.
5 Validation of AI-generated MOFs with molecular dynamics simulations
We now examine the stability and porous properties of the 6,020 MOFs described in the previous section by using molecular dynamics (MD) simulations. We use the LAMMPS [53] code, version 28 Mar 2023 Update 1, distributed from the conda-forge channel, to perform MD simulations that relax the AI-generated MOF structures. An adapted version of cif2LAMMPS [54] is used to generate LAMMPS input files automatically. For each MOF, a 2×2×2 supercell structure is generated, and an NPT simulation is performed, i.e., we use an isothermal-isobaric ensemble with a constant particle number $N$, a pressure $p$ fluctuating around an equilibrium value, $\langle p\rangle=1\textrm{atm}$, and a temperature $T$ fluctuating around an equilibrium value $\langle T\rangle=300\textrm{K}$. The cell lengths and angles of the triclinic simulation box are allowed to relax.
We run the NPT simulations for 200,000 steps with a timestep of 1 femtosecond. The relative changes of structure density before and after the simulation are inspected to evaluate how well the porous characteristics are maintained throughout the simulation. This process produces 102 MOFs whose structures stayed intact during the MD simulation. Next, we use a $<$1% change in density as the threshold to determine whether a MOF structure is stable or not. Among the 102 MOF candidates, 18 have change in density below the 1% threshold. Figure 9 presents the top five AI-generated, stable MOF structures that passed the threshold for structure determination—see Table 4 for combinations of their building blocks and Figure 10 for 3D visualization of their crystal structures.
Figure 9: Top five AI-generated MOFs with structure density changed less than 1% during molecular dynamics simulations. Their CO2 capacities are higher than 96.9% of hMOF structures in the hMOF database.
Figure 10: 3D visualization of the crystal structures of top five AI-generated, stable MOF candidates that retained their porous properties throughout molecular dynamics simulations. Atoms are presented using the color code: carbon in grey, nitrogen in dark blue, fluorine in cyan, zinc in purple, hydrogen in white, and lithium in green.
6 Conclusion
We have introduced GHP-MOFassemble, an AI-driven discovery framework for high-throughput generation of novel high-performing MOF structures for carbon capture. The GHP-MOFassemble framework employs a diffusion model, DiffLinker, to generate new, unique, and chemically diverse MOF linkers, which it then assembles with one of three pre-selected metal nodes into MOFs in the pcu topology.
Out of 120,000 MOF structures generated by GHP-MOFassemble, 6,020 were predicted to be high-performing MOFs with CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of the hMOF dataset. We relaxed each of these 6,020 MOFs in a 200-picosecond NPT molecular dynamics simulation with the UFF4MOF force field [55, 56], and found that 102 structures did not collapse after molecular dynamics simulations, among which 18 were identified to be stable structures with less than 1% change in density.
We have deployed and extensively tested GHP-MOFassemble on computers at the Argonne Leadership Computing Facility and at the National Center for Supercomputing Applications, with the intent of providing scalable and computationally efficient AI tools to accelerate the modeling and discovery of novel MOF structures. The tools introduced in this work may be readily fine-tuned and adapted to other available datasets beyond hMOF to enable accelerated design and discovery of novel MOF structures for carbon capture at industrial scale.
7 Acknowledgments
This work was supported by Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, provided by the Director, Office of Science, of the U.S. Department of Energy under Contract No. DE-AC02-06CH11357, and by the Braid project of the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under contract number DE-AC02-06CH11357. The work used resources of the Argonne Leadership Computing Facility, a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357. EAH and IF acknowledge support from National Science Foundation (NSF) award OAC-2209892. SC and XY acknowledge partial support from NSF Future of Manufacturing Research Grant 2037026. This research also used the Delta advanced computing and data resources, which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois at Urbana-Champaign and its National Center for Supercomputing Applications.
Code availability
The code and data are available upon reasonable request from the authors.
ORCID IDs
Hyun Park $0$-0001-5550-5610
Xiaoli Yan $0$-0002-6512-2338
Ruijie Zhu $0$-0001-9316-7245
Eliu Huerta $0$-0002-9682-3604
Santanu Chaudhuri $0$-0002-4328-2947
Ian Foster $0$-0003-2129-5269
Emad Tajkhorshid $0$-0001-8434-1010
Appendix A Dummy atom identification
Figure 11 illustrates the workflow used to identify the dummy atoms and to remove redundant atoms for linkers containing carboxyl groups (left panel) or heterocyclic rings (right panel).
Figure 11: Left panel, identification of dummy atoms for linkers containing carboxyl groups. The dummy atoms are found by substituting the carbon atoms in the carboxyl groups. The remaining oxygen and hydrogen atoms in the carboxyl groups are removed. Right panel, identification of dummy atoms for linkers containing heterocyclic rings. The dummy atoms are found at nitrogen-metal bond distance from the terminal nitrogen atoms along the vectors pointing from the opposing carbon atoms to nitrogen atoms.
Appendix B MOF catenation via site translation
Catenated MOFs were generated by using the site translation method, which is achieved by displacing the reference lattice along the diagonal line of the unit cell. For all four catenation levels, the amount of relative lattice displacement is given in Table 5. The numbers are fractional displacements relative to the unit cell diagonal.
Table 5: Amount of lattice displacement for catenated MOFs
catenation lattice1 lattice2 lattice3 lattice4
cat0 (0,0,0) / / /
cat1 (0,0,0) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) / /
cat2 (0,0,0) ($\frac{1}{3}$,$\frac{1}{3}$,$\frac{1}{3}$) ($\frac{2}{3}$,$\frac{2}{3}$,$\frac{2}{3}$) /
cat3 (0,0,0) ($\frac{1}{4}$,$\frac{1}{4}$,$\frac{1}{4}$) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) ($\frac{3}{4}$,$\frac{3}{4}$,$\frac{3}{4}$)
Appendix C Linker evaluation
Recall from Section 4.2 that we use the synthetic accessible score (SAscore) and synthetic complexity score (SCscore) metrics to confirm that all linkers are synthesizable. Table 2 provides additional data on the generated linkers. We also provide results for the validity and uniqueness metrics, and the internal diversity scores IntDiv1 and IntDiv2 [48]. The internal diversity scores were calculated based on the Tanimoto distances among all pairs of molecules, which are obtained by calculating the normalized Jaccard score of Morgan fingerprint bit vector between all pairs of generated linkers. This yields similarities between a specific linker and the rest of the linkers in the linker pool, which are averaged out. Therefore, we end up with a metric showing each linker’s similarity compared with the population of generated linkers. We used the MOSES [48] framework to compute these two scores using the relations:
$\textrm{IntDiv}_{1}(G)=1-\frac{1}{|G|^{2}}\sum_{m_{1},m_{2}\in G}T_{d}(m_{1},m_{2})\,,$ (7)
$\textrm{IntDiv}_{2}(G)=1-\sqrt{\frac{1}{|G|^{2}}\sum_{m_{1},m_{2}\in G}T_{d}(m_{1},m_{2})^{2}}\,,$ (8)
where $G$ is the generated set of molecules, $|G|$ is the size of molecule set, $(m_{1},m_{2})$ represents a pair of molecules in the molecule set. $T_{d}$ is the Tanimoto distance, which relates to the Tanimoto similarity ($T_{s}$) by:
$T_{d}(m_{1},m_{2})=1-T_{s}(m_{1},m_{2})$ (9)
Appendix D Statistics and distribution of the standard deviation of the ensemble of pre-trained models
Table 6 shows the R2 score, mean absolute error (MAE), and root mean squared error (RMSE) of the three pre-trained models used for creating the ensemble model. Figure 12 shows the distribution of the standard deviations of ensemble model predictions.The threshold 0.2 mmol/g serves as a criterion for assessing whether the predictions made by the three pre-trained models are in agreement. The ensemble model predictions with standard deviation above the threshold were discarded.
Table 6: Statistics of the three pre-trained models
Model R2 MAE RMSE
Model1 0.932 0.098 0.171
Model2 0.937 0.100 0.170
Model3 0.936 0.099 0.170
Figure 12: Distribution of the standard deviation of ensemble model predictions. The ensemble model predictions with standard deviation of less than 0.2 mmol/g consist of around 96% of the test set. This shows that our three independently trained models are in great agreement for most of the unobserved test set, showing robustness of our model. The remaining 4% with larger than 0.2 mmol/g standard deviation implies that these are the data points which may be difficult to predict due to errors in target property or extra information other than atomic species and periodic neighbor is necessary for accurate prediction.
Appendix E Linker similarity for high-performing MOF structures
To measure the similarity between the generated linkers and hMOF linkers in high-performing MOF structures, we show in Figure 13 the distribution of maximum Tanimoto similarity between the two linker sets. For each unique linker in the predicted high-performing MOF structures generated by the GHP-MOFassemble framework, we calculate its Tanimono similarity with all of the unique linkers in the hMOF structures. The maximum value of the Tanimoto similarities gives a quantitative measure of how different the generated linkers are from the hMOF linkers.
Figure 13: Distribution of maximum Tanimono similarity between generated and hMOF linkers in high-performing MOF structures. The peak around 0.3 to 0.4 indicates that most generated linkers are just 30–40% similar to those in hMOF–showing that we are generating novel linkers not present in hMOF structures. On the other hand, the trailing heavy right tail above 0.4 Tanimoto similarity indicates that we are also able to generate linkers that are structurally similar to those present in hMOF, showing that GHP-MOFassemble enables generation of a diverse set of novel linkers.
Appendix F Regression model training details
We independently trained three modified versions of the CGCNN model for 5,000 epochs with a batch size of 160, and with optimizer, learning rate, and weight decay of torch_adam, 1e-4, and 2e-5, respectively.
Appendix G Linker statistics
Table 7: Statistical summary of the number of linkers after each step categorized by the corresponding MOF types. PW and TM stand for paddlewheel and tetramer, respectively. Element filter means that linkers with S, Br and I elements are removed.
Method Total Cu PW-pcu Zn PW-pcu Zn TM-pcu
DiffLinker 64,800 21,600 19,440 23,760
Hydrogen addition 56,257 18,979 17,126 20,152
Dummy atom identification 16,162 4,964 4,450 6,748
Element filter 12,305 3,702 3,441 5,162

### v2

@@
-GHP-MOFassemble: Diffusion modeling, high throughput screening, and molecular dynamics for rational discovery of novel metal-organic frameworks for carbon capture at scale
+A generative artificial intelligence framework based on a molecular diffusion model for the design of metal–organic frameworks for carbon capture
 Abstract
-We introduce GHP-MOFassemble, a Generative artificial intelligence (AI), High Performance framework to accelerate the rational design of metal-organic frameworks (MOFs) with high CO2 capacity and synthesizable linkers.
-Our framework combines a diffusion model, a class of generative AI, to generate novel linkers that are assembled with one of three pre-selected nodes into MOFs in a primitive cubic (pcu) topology.
-The CO2 capacities of these AI-generated MOFs are predicted using a modified version of the crystal graph convolutional neural network model.
-We then use the LAMMPS code to perform molecular dynamics simulations to relax the AI-generated MOF structures, and identify those that converge to stable structures, and maintain their porous properties throughout the simulations.
-Among 120,000 pcu MOF candidates generated by the GHP-MOFassemble framework, with three distinct metal nodes (Cu paddlewheel, Zn paddlewheel, Zn tetramer), a total of 102 structures completed molecular dynamics simulations at 1 bar with predicted CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of hMOFs in the hypothetical MOF (hMOF) dataset in the MOFX-DB database.
-Among these candidates, 18 have change in density lower than 1% during molecular dynamics simulations, indicating their stability.
-We also found that the top five GHP-MOFassemble’s MOF structures have CO2 capacities higher than 96.9% of hMOF structures.
-This new approach combines generative AI, graph modeling, large-scale molecular dynamics simulations, and extreme scale computing to open up new pathways for the accelerated discovery of novel MOF structures at scale.
-  • July 13, 2023
-Keywords: diffusion model, pre-trained regression model, metal-organic framework, carbon capture
-1 Introduction
+Metal–organic frameworks (MOFs) exhibit great promise for CO2 capture.
+However, finding the best performing materials poses computational and experimental grand challenges in view of the vast chemical space of potential building blocks.
+Here, we introduce GHP-MOFassemble, a generative artificial intelligence (AI), high performance framework for the rational and accelerated design of MOFs with high CO2 adsorption capacity and synthesizable linkers.
+GHP-MOFassemble generates novel linkers, assembled with one of three pre-selected metal nodes (Cu paddlewheel, Zn paddlewheel, Zn tetramer) into MOFs in a primitive cubic topology.
+GHP-MOFassemble screens and validates AI-generated MOFs for uniqueness, synthesizability, structural validity, uses molecular dynamics simulations to study their stability and chemical consistency, and crystal graph neural networks and Grand Canonical Monte Carlo simulations to quantify their CO2 adsorption capacities.
+We present the top six AI-generated MOFs with CO2 capacities greater than 2m $\textrm{m\,mol}\,\textrm{g}^{-1}$, i.e., higher than 96.9% of structures in the hypothetical MOF dataset.
+Introduction
 Metal-organic frameworks (MOFs) have garnered much research interest in recent years due to their diverse industrial applications, including gas adsorption and storage [1], catalysis [2], and drug delivery [3].
@@
 MOFs have been shown to exhibit superior chemical and physical CO2 adsorption properties.
-They can be recycled in appropriate operating environments for varying numbers of times before undergoing significant structural degradation.
+In appropriate operating environments, they can be recycled, for varying numbers of times, before undergoing significant structural degradation.
 However, their industrial applications have not yet reached full potential due to stability issues, such as poor long-term recyclability [5], and high moisture sensitivity [6].
 For instance, the presence of moisture in adsorption gas impairs a MOF’s CO2 capture performance, which may be attributed to MOFs having stronger affinity toward water molecules than to CO2 molecules [7].
-Other investigations of MOF stability issues [8, 9, 10] have shown [11, 12] that the gas adsorption properties of MOFs can be enhanced by tuning the building blocks used.
-Yet progress has been difficult due to the enormous chemical space of building blocks that makes exhaustive search with traditional experimental methods impractical [13].
-Here we propose GHP-MOFassemble, a novel high-throughput computational framework to accelerate the discovery of MOF structures with high CO2 capacities and synthesizable linkers.
-To demonstrate the utility of our framework, we apply it to the rational design of MOFs with pcu topology and three types of inorganic nodes: Cu paddlewheel, Zn paddlewheel, and Zn tetramer.
-The key component of our generative framework is a diffusion model, Difflinker [14], which enables efficient generation of new linker molecules from a set of molecular fragments.
-Previous attempts to discover useful MOFs have proceeded via high-throughput screening of existing datasets [15, 16, 17], an approach that necessarily limits the search to known MOFs.
-Our GHP-MOFassemble framework instead probes the MOF design space by employing a molecular generative diffusion model (DiffLinker [14]) to generate chemically diverse and unique MOF linkers, which it assembles with pre-selected metal nodes to form novel MOF structures.
-It then screens those structures with a pre-trained regression model, a modified version of CGCNN [18], to identify high-performing MOF candidates for carbon capture.
-2 Related Work
-Previous efforts in the search for new MOF structures with exceptional gas adsorption properties include high-throughput screening and generative modeling.
-High-throughput screening approaches comprise both database search and machine learning (ML)–assisted screening.
+Other investigations of MOF stability issues [8, 9, 10, 11, 12] have shown that the gas adsorption properties of MOFs can be enhanced by tuning the building blocks used.
+Yet progress has been difficult due to the enormous chemical space of building blocks, which makes exhaustive search with traditional experimental methods impractical [13].
+Related Work
+Previous efforts in the search for new MOF structures with exceptional gas adsorption properties include:
 Database search methods.
-These methods apply filters to identify attractive candidates in large databases of MOF structures plus calculated properties obtained by using molecular simulations.
-For example, Li et al.
-[16] search the computation-ready, experimental MOF database (CoRE DB) for MOFs with high $\textrm{CO}_{2}$ uptake when moisture is present, and Altintas et al.
-[19] search both CoRE DB and the Cambridge Structural Database non-disordered MOF subset for MOFs with high $\textrm{CH}_{4}/\textrm{H}_{2}$ selectivity.
-ML–assisted screening.
-Database search methods require expensive molecular simulation calculations for every target molecule.
-ML-assisted screening avoids this difficulty by using a regression model trained on a smaller training dataset to predict target properties for a large number of new test structures.
+These methods apply filters to identify optimal candidates in large databases of MOF structures plus calculated properties obtained with molecular simulations.
+For example, Ref.
+[14] shows how to search for MOFs with high $\textrm{CO}_{2}$ uptake when moisture is present using the experimental MOF database (CoRE DB), while Ref.
+[15] shows how to search for MOFs with high $\textrm{CH}_{4}/\textrm{H}_{2}$ selectivity in both CoRE DB and the Cambridge Structural Database non-disordered MOF subset.
+Machine Learning (ML)-assisted screening.
+Building large MOF databases require expensive molecular simulation calculations for every target structure.
+ML-assisted screening avoids this difficulty by using a regression model trained on a smaller training set to predict target properties of a large number of new test structures.
 This approach has been widely applied to gas adsorption and separation property predictions of MOFs, including CO2/H2 separation [16] and CH4 adsorption [17].
-ML-assisted methods commonly make use of both geometrical features (e.g., dominant pore size, void fraction [22]) and chemical features (e.g., atom type, electronegativity [23]).
-A compound feature, the atomic property weighted radial distribution function (AP-RDF) [24] has been shown to improve regression model performance in finding MOF structures with high CO2 capacity [25].
-This feature is constructed by using both chemistry and local geometry of atomic sites in MOF structures.
-As an alternative to the use of hand-engineered machine learning features, neural networks have also been applied to MOF research.
-In a recent study, the Atomistic Line Graph Neural Network (ALIGNN), trained on the hMOF dataset [26], was applied to search CoRE DB for high-performing MOF candidates for carbon capture [27].
+ML-assisted methods use of geometrical features (e.g., dominant pore size, void fraction [18]) and chemical features (e.g., atom type, electronegativity [19]).
+A compound feature, the atomic property weighted radial distribution function [20] has been shown to improve regression model performance in finding MOF structures with high CO2 capacity [21].
+This feature is constructed by using both local geometry and chemistry of atomic sites in MOF structures.
+As an alternative to the use of hand-engineered ML features, neural networks have been applied to MOF research.
+In Ref.
+[22], the Atomistic Line Graph Neural Network, trained on the hMOF dataset, was applied to search CoRE DB for high-performing MOFs for carbon capture [23].
 Generative modeling.
-In the generative modeling approach, candidate molecules are not drawn from a database but are generated de novo via methods that produce novel compounds that are expected to have a desired set of chemical features.
-One notable example is the recently proposed Supramolecular Variational Autoencoder (SmVAE[28], which uses a semantically constrained graph-based canonical code (RFcode) to encode MOF building blocks.
+Candidates are not drawn from a database but are generated de novo via methods that produce novel compounds that have a desired set of chemical features.
+For instance, the Supramolecular Variational Autoencoder [24] uses a semantically constrained graph-based canonical code to encode MOF building blocks.
 The variational autoencoder framework structure allows this model to interpolate between existing MOF structures and enables isoreticular optimization of MOF structures toward higher CO2 capacity and CO2/N2 selectivity.
 This work.
-Proposed generative models include not only variational autoencoders (VAEs) but also generative adversarial networks (GANs), normalizing flows, autoregressive models, and diffusion models [29].
-In this work, we adopt a diffusion model named DiffLinker to generate novel MOF linkers.
-Diffusion models use a probability distribution and Markovian properties to generate new data via a denoising step.
+Proposed generative models include variational autoencoders, generative adversarial networks, normalizing flows, autoregressive models, and diffusion models [25].
+Here, we adopt a diffusion model named DiffLinker to generate novel MOF linkers.
+Diffusion models use a probability distribution and Markovian properties to generate new data via forward diffusion and backward denoising steps.
 First, Gaussian noise is added to the input samples to yield noisy data.
@@
 The trained network is then used to reversely transform (i.e., denoise) the noisy data back into target samples that resemble molecules from the training data distribution.
-This method has been widely applied to drug discovery to speed up the design of new ligands and ligand-protein docking configurations [30, 31, 32, 33].
+This method has been widely applied to drug discovery to speed up the design of new ligands and ligand-protein complexes [26, 27, 28, 29].
 Two broad categories of molecular representation schemes have been used in diffusion model-based ligand generators: molecular graphs and 3D coordinates.
 In the former case (Digress/Congress [30] is an example), atoms are represented as nodes and bonds as edges.
-In the latter case, the 3D atomic coordinates of molecules are generated directly, as in DiffLinker [14] and E(3) equivariant diffusion model (EDM[35].
-Given the success of diffusion models in drug discovery [32, 30, 33, 36, 37], in Section 3.1 we demonstrate how to transfer the idea to the iso-reticular design of MOF structures by varying MOF linkers while fixing node and topology.
-The diffusion model is used specifically for MOF linker design, as we describe in Section 3.2.
-3 Methods
+In the latter case, the 3D atomic coordinates of molecules are generated directly, as in DiffLinker [31] and E(3) equivariant diffusion model (EDM) [32].
+Given the success of diffusion models in drug discovery [28, 26, 29, 33, 34], we demonstrate how to transfer the idea to the iso-reticular design of MOF structures by varying MOF linkers while fixing node and topology.
+The diffusion model is used specifically for MOF linker design.
+Our proposed approach, GHP-MOFassemble, is a novel high-throughput computational framework to accelerate the discovery of MOF structures with high CO2 capacities and synthesizable linkers.
+While previous attempts to discover useful MOFs via computational methods have proceeded via high-throughput screening of existing datasets [35, 14, 36], an approach that necessarily limits the search to known MOFs.
+In contrast, GHP-MOFassemble probes the MOF design space by employing a molecular generative diffusion model, DiffLinker [31], to generate chemically diverse and unique MOF linkers from a set of molecular fragments, which it assembles with pre-selected metal nodes to form novel MOF structures.
+It then screens those structures with a pre-trained regression model, a modified version of Crystal Graph Convolutional Neural Network (CGCNN) [37], to identify high-performing MOF candidates for carbon capture.
+In addition, we apply molecular dynamics (MD) and grand canonical Monte Carlo (GCMC) to further down-select stable AI-generated MOFs and compute more credible CO2 adsorption capabilities.
+We demonstrate the utility of our framework by applying it to the rational design of MOFs with pcu topology and three types of inorganic nodes: Cu paddlewheel, Zn paddlewheel, and Zn tetramer.
+Results
+Here we describe the key components of GHP-MOFassemble, and present a new set of AI-generated MOF structures with high CO2 capacities.
+A detailed description of the approaches used to obtain these findings is provided in Methods.
+Analysis of the hMOF dataset
+The three most frequent node-topology pairs in the hMOF dataset, accounting for around 74% of its MOFs, are the Cu paddlewheel-pcu, Zn paddlewheel-pcu, and Zn tetramer-pcu, which amount to 102,117 hMOF structures (29,714 + 28,529 + 43,874 from first column of Table 1).
+Out of these 102,117 structures, we only used those with correctly parsed MOFids and valid Simplified Molecular-Input Line-Entry System (SMILES), i.e., 78,238 MOFs.
+Table 1: Most frequent node-topology pairs in the hMOF dataset. Properties of the hMOF dataset. PW, TM, and HP-MOF stand for paddlewheel, tetramer, and high-performing MOF, respectively.
+node-topology total MOFs HP-MOFs HP-MOFs with unique linker unique molecular
+three linkers SMILES fragment conformers
+Cu PW-pcu 29,714 1,458 1,016 3,330 180
+Zn PW-pcu 28,529 1,314 834 3,221 162
+Zn TM-pcu 43,874 2,129 1,388 3,265 198
+Linker generation and evaluation
+For all three MOF types, we start with 540 (i.e., = 180 + 162 + 198 from Table 1 last column) unique molecular fragments extracted from high-performing hMOF structures, and use DiffLinker to generate new MOF linkers, with the number of sampled connection atoms ranging from five to 10 (total six connection atom types).
+For each linker, sampling is performed 20 times.
+Therefore, the resulting candidate pool contains a total of 64,800 linkers, i.e., 540 fragments $\times$ 20 independent sampling steps $\times$ six unique number of connection atoms.
+Since these linkers only contain heavy atoms, to generate all-atom linker molecules, we apply openbabel to add hydrogen atoms, which results in 56,257 linkers after removing linkers with erroneous hydrogen assignments.
+Next, dummy atom identification is performed to generate information that enables assembly with metal nodes.
+A total of 16,162 linkers with dummy atoms are generated.
+These linkers are then passed through the element filter, which removes linkers that contain S, Br, and I, further reducing the number to 12,305 linkers.
+The number of linkers in each step are summarized in the Total column of Table 2.
+Table 2: Number of linkers at each step of the MOF assembly process.
+Statistical summary of the number of linkers after each step categorized by the corresponding MOF types.
+PW and TM stand for paddlewheel and tetramer, respectively.
+Element filter means that linkers with S, Br and I elements are removed.
+Method Total Cu PW-pcu Zn PW-pcu Zn TM-pcu
+DiffLinker 64,800 21,600 19,440 23,760
+Hydrogen addition 56,257 18,979 17,126 20,152
+Dummy atom identification 16,162 4,964 4,450 6,748
+Element filter 12,305 3,702 3,441 5,162
+MOF assembly
+We generate new MOF structures by assembling three randomly selected DiffLinker-generated linkers with one of the three most frequent nodes in the hMOF dataset.
+Random sampling of 12,305 linkers ensures that the selected linkers cover a large design space.
+For each node-linker-topology combination, we considered four levels of catenation (cat0, cat1, cat2, cat3).
+We generated a total of 120,000 MOFs as follows: we have four catenation levels and three node candidates.
+The random sampling of 10,000 linkers for each catenation level-node candidate pair generates 120,000 total MOFs of different catenation-node-linker combinations.
+MOF geometry examination
+Inter-atomic distance check
+We used Pymatgen to read each MOF structure’s CIF file and to extract the pairwise distance matrix.
+Diagonal entries of the distance matrix are discarded as they signify the distance between an atom and itself.
+The off-diagonal values are tested for minimum inter-atomic distance threshold.
+The inter-atomic distance threshold is predetermined by an experimental database, OChemDb [38].
+If a MOF has at least one entry of lower than threshold inter-atomic distance, the MOF is discarded.
+78,796 of the 120,000 assembled MOFs passed this test.
+Pre-simulation Check
+We used the open source library, cif2lammps [39], to automatically assign the Universal Force Field for Metal–Organic Frameworks (UFF4MOF) [40, 41] to MOFs and generate LAMMPS [42] input files.
+This step ensures that all atomic structures and bonding appearing in each MOF structure are chemically valid within the scope of UFF4MOF.
+For example, if a Zn atom is bonded to a C atom, or if an O atom is bonded to four other atoms, etc., these structures will be identified as invalid and will be discarded.
+Of the 78,796 MOFs that passed geometry and inter-atomic distance checks, 18,770 passed this pre-simulation check, and thus had LAMMPS input files generated.
+Regression model for MOF $\textrm{CO}_{2}$ capacity prediction
+To reduce the number of LAMMPS simulations, a screening step of the MOFs’ adsorption performance is conducted on the 18,770 MOFs described above using a modified version of the CGCNN model that we introduced in Ref.
+[37].
+Here, we used MOF structures from the hMOF dataset as well as their CO2 capacities at 0.1 bar as input data.
+We split the hMOF dataset into three independent sets: 80% for training, 10% for validation, and 10% for testing.
+Using this data split, we trained three CGCNN models using random initialization of weights for each of them.
+When we use this model ensemble to infer the CO2 capacities of newly AI-generated MOF structures, we take the average of the predictions made by the three independent models as the predicted CO2 adsorption capacity.
+Out of the 18,770 AI-generated MOF structures, a total of 364 were predicted to be high-performing MOFs with CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ at 0.1 bar.
+Specifically, we used ensemble model prediction mean plus standard deviation as CO2 adsorption capacity value to be higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ as the threshold.
+The standard deviation consideration is to assure we account for statistical errors of adsorption predictions by our ensemble model.
+Categorization of the predicted high-performing MOFs by node-topology pair and catenation level indicates that for all three node types most of the high-performing MOFs are cat2 and cat3.
+Structural validation of AI-generated MOFs with molecular dynamics simulations
+We now examine the stability and porous properties of the 364 AI-generated MOFs described in the previous section using MD simulations with LAMMPS [42].
+For each MOF, a 2×2×2 supercell structure is equilibrated under a triclinic isothermal-isobaric ensemble (i.e., NPT) at $\langle p\rangle=1\ \textrm{atm}$ and $\langle T\rangle=300\textrm{K}$, such that the cell lengths and angles of the MOF structure can be equilibrated.
+The NPT simulations are run for 400,000 steps with step size of 0.5 femtoseconds.
+The relative changes of the lattice parameters before and after the simulation are inspected to evaluate how well the MOF structure is maintained throughout the equilibrium MD simulation.
+The MOFs with higher than 5% changes in any of the lattice parameters ($a$, $b$, $c$, $\alpha$, $\beta$, $\gamma$) are discarded (first three are lattice vector lengths and the latter three are lattice angles) between the initial state and the equilibrated state.
+The MOFs with higher than 5% changes in any of the lattice parameters ($a$, $b$, $c$, $\alpha$, $\beta$, $\gamma$) are discarded.
+This process produces 102 MOFs that satisfied the previously defined criteria, i.e., $<5\%$ in lattice parameter change [40].
+Property validation of AI-generated MOFs with grand canonical Monte Carlo (GCMC) simulations
+These 102 MOFs are further examined with GCMC simulations to calculate their CO2 adsorption capacity at a pressure and temperature of 0.1 bar and 300 K, respectively.
+RASPA [43] is used to conduct these GCMC CO2 adsorption simulations.
+As demonstrated in Figure 1, 6 MOFs candidates were found to have CO2 adsorption capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ by the GCMC simulations.
+The 3D structure of these six high performing, AI-generated MOFs is presented in Figure 2.
+Figure 1: CO2 adsorption values at 0.1 bar and 300K.
+CO2 adsorption capacity values at a pressure and temperature of 0.1 bar and 300 K, respectively, for the top six AI-generated MOFs’ according to grand canonical Monte Carlo (GCMC) simulations and our modified crystal graph convolutional neural network (CGCNN) model.
+Figure 2: Visualization of the crystal structure of AI-generated MOFs.
+a–f Crystal structure of the top six AI-generated MOFs.
+The color code used to represent atoms is: carbon in grey, nitrogen in dark blue, fluorine in cyan, zinc in purple, hydrogen in white, and lithium in green.
+Discussion
+GHP-MOFassemble combines novel generative and graph AI applications, as well as a comprehensive screening workflow that combines various modeling methods with increasing levels of chemistry awareness and increasing levels of computational cost.
+When we deployed and optimized GHP-MOFassemble on the Delta supercomputer, unless indicated otherwise, we found that:
+  • We AI-assembled 120,000 MOFs within 33 minutes using multiprocessing on 28 cores in the ThetaKNL supercomputer at the Argonne Leadership Computing Facility (ALCF).
+  • We screened these 120,000 MOFs in 40 minutes using multiprocessing on 128 cores, and identified 78,796 MOFs with valid bond lengths.
+  • We further screened these 78,796 MOFs and identified 18,770 MOFs with valid chemistry, using UFF4MOF as a reference, within 205 minutes using multiprocessing on 128 cores.
+  • We then used our AI ensemble of CGCNN models to estimate the CO2 capacity of these 18,770 MOFs, and identified 364 high-performing, AI-generated MOFs with CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ at 0.1 bar.
+AI inference was completed in 50 minutes using one NVIDIA A40 GPU.
+In brief, from assembly to selection of high-performing MOFs, GHP-MOFassemble completes the analysis within 5 hours and 7 minutes.
+Once we have selected 364 AI-generated, high-performing MOFs, we carry out the most compute-intensive part of the analysis, namely:
+  • We used the LAMMPS code to equilibrate 364 MOFs with 200-picosecond NPT MD simulations with the UFF4MOF force field.
+Each of these simulations is completed within 11 minutes, on average, using between 6 to 14 MPI processes, depending on the number of atoms in the MOF.
+We identified 102 stable MOFs whose lattice parameters changed less than 5% throughout the MD simulations.
+  • Finally, each of these 102 MOFs were further examined with GCMC simulations, and a total of six MOF candidates were found to have CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$, which corresponds to the top 5% of the hMOF dataset.
+Each of these 102 GCMC simulations is completed within six hours, on average, using one CPU core.
+Recent studies [44, 45, 46, 47] indicate that several functional groups play an important role in determining MOF CO2 capacity, including carboxylic acid (-COOH), primary amine (-NH2), hydroxyl (-OH), and nitrile (-CN).
+These functional groups affect MOF CO2 capacity due to their interaction with CO2 molecules through charge redistribution [46], hydrogen bond interaction [45], and electrostatic interaction [45].
+Thus, we analyzed the functional groups of linkers in the 364 AI-generated high-performing MOF structures and compared them with high-performing hMOF structures.
+The results of this analysis are presented in Figure 3.
+We observe that linkers in high-performing hMOF structures have a higher proportion of carboxylic groups, whereas hydroxyl groups appear more often in the predicted high-performing MOFs generated by our framework.
+Moreover, the percentages of primary amine and nitrile in linkers are similar in both cases.
+Note that some substructures appear more often than others.
+For example, ring structures appear in most of the linkers in the top candidates, and many of the DiffLinker-generated molecular substructures (in between the terminal fragments) also contain ring structures.
+The frequent occurrence of rings in linkers of high-performing MOFs may be due to the presence of CO2-ring interaction, as reported in Ref.
+[48].
+In addition, rings in the linkers may increase MOFs’ structural rigidity due to less rotatable bonds.
+In summary, the entire discovery process from generative AI MOF assembly to detailed GCMC simulations of top AI-predicted MOFs can be completed within 12 hours using distributed computing in modern supercomputing platforms.
+The entire workflow may be further accelerated by scaling up any of the subprocesses using more cores or by distributing AI inference over more GPUs.
+Furthermore, the various pieces of GHP-MOFassemble may be assembled to create a standalone workflow that combines MD, density functional theory, and GCMC simulations to expose our generative AI model to a larger set of chemically diverse, high performing MOFs.
+Through online learning methods one may continually guide generative AI until it consistently assembles high performing MOFs.
+A study of this nature will be pursued in future work.
+Figure 3: Functional groups in high-performing MOFs. Comparison of the proportion of selected functional group that appear in high-performing AI-generated MOFs and MOFs in the hMOF dataset.
+Methods
 Our GHP-MOFassemble framework has three components:
   • Decompose.
-We use a molecular fragmentation algorithm to decompose the MOF linkers found in high-performing MOFs within the hMOF dataset [26]—an open source dataset that provides, for each of 137,652 hypothetical MOF structures, its MOFid, MOFkey, geometric features, and isotherm data for six adsorption gases (CO2, N2, CH4, H2, Kr, Xe) at 0.01, 0.05, 0.1, 0.5, and 2.5 bar—covering the pressure ranges of cyclic adsorption gas separation processes in industrial applications.
-The adsorption properties of gases provided in this dataset were calculated using Grand Canonical Monte Carlo (GCMC) calculations [26], as described in Wilmer et al.
+We use a molecular fragmentation algorithm to decompose the MOF linkers found in high-performing MOF structures within the hMOF dataset [22]—an open source dataset that provides, for each of 137,652 hypothetical MOF structures, and corresponding MOFid, MOFkey, geometric features, and isotherm data for six adsorption gases (CO2, N2, CH4, H2, Kr, Xe) at 0.01, 0.05, 0.1, 0.5, and 2.5 bar—covering the pressure ranges of cyclic adsorption gas separation processes in industrial applications.
+The adsorption properties of gases provided in this dataset were calculated using GCMC calculations [22], as described in Ref.
 [49].
@@
 We use a diffusion model to generate new MOF linkers.
-We then screen the generated linkers by removing linkers with S, Br and I elements (we call this step “element filter” in later text), and evaluate their quality using five different scores to quantify their synthesizability, validity, uniqueness, and internal diversity.
+We then screen the AI-generated linkers by removing linkers with S, Br and I elements (we call this step “element filter”), and evaluate their quality using five different scores to quantify their synthesizability (SAScore and SCScore), validity, uniqueness, and internal diversity.
 Linkers that passed the element filter are then assembled with one of three pre-selected nodes into new MOF structures in the pcu topology.
-  • Predict. We use a pre-trained regression model to predict the CO2 capacities of the newly generated MOF structures.
-In the following sections we describe each of these components in detail.
-3.1 Decomposing MOF linkers into molecular fragments
+  • Screen and Predict.
+We check for inter-atomic distances with pre-simulation checks to ensure we filter poor AI-generated MOF structures from previous steps.
+We then use a pre-trained regression model to predict the CO2 capacities of the newly generated MOF structures.
+Lastly, we perform MD and GCMC simulations to obtain stable and high-performing MOF structures.
+In the following sections we describe the datasets we have used, and each of the GHP-MOFassemble components in detail.
+The hMOF dataset
+We selected the most common topologies of the hMOF dataset for this work, i.e., Cu paddlewheel-pcu, Zn paddlewheel-pcu, and Zn tetramer-pcu, i.e., 78,238 MOFs with correctly parsed MOFids and valid SMILES.
+We gather the numbers of molecular fragments in Table 1, produced through the Fragment step of our GHP-MOFassemble framework, described below.
+In Table 1 high-performing MOFs (second column) with three parsed linkers are selected (third column), and their unique linkers are parsed by using the MMPA (Matched Molecular Pairs Algorithm) algorithm.
+The cumulative distribution functions of the CO2 capacities of these three types of MOFs are shown in the right panel of Figure 4.
+Figure 4: Properties of the hMOF dataset. a Depictions of the most frequent node-topology pairs in hMOF structures. b Cumulative distribution functions of their 0.1 bar CO2 capacities.
+In Table 1, the number of output molecular fragment conformers (last column) is much less than the number of unique linker SMILES (second to last column) for two reasons.
+First, around 56% of linkers did not pass the valency check, which may be due to the intrinsic limitations in the parsing of SMILES of MOF linkers.
+Second, around 90% of linkers that passed the valency check share similar molecular fragments, and thus many duplicated molecular fragments exist.
+The successfully parsed unique molecular fragment conformers are subsequently used for linker generation.
+Figure 5 shows the empirical cumulative distribution functions of the CO2 capacities of hMOF structures at different catenation levels (i.e., MOFs with interpenetrated lattices).
+Therein, we observe that a higher percentage of catenated MOFs (cat1, cat2 and cat3) are high-performing, as compared to the uncatenated MOFs (cat0).
+This result confirms that catenation is an important factor when designing new MOF structures with high CO2 capacity.
+This observation is consistent with other studies in the literature [50, 51], which indicate that even though catenation reduces pore size and surface areas, catenated MOFs generally have higher CO2/H2 selectivities because MOF-CO2 interactions are enhanced as a result of the strong confinement of CO2 with a much lower adsorption surface.
+Thus, the results presented in Figure 5 using the hMOF dataset indicate that CO2 working capacities of catenated MOFs are higher than their noncatenated counterparts.
+As we discuss below, we found a similar pattern in AI-generated MOFs.
+Catenated MOFs were generated by using site translation method, which is achieved by displacing the reference lattice along the diagonal line of the unit cell.
+For all four catenation levels, the amount of relative lattice displacement is given in Table 3.
+The numbers are fractional displacements relative to the unit cell diagonal.
+Figure 5: CO2 capacities of hMOF structures at different catenation levels.
+Empirical cumulative distribution functions of MOFs in the hMOF dataset at 0.1 bar at different catenation levels.
+The x axis is capped at 4 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ to preserve details.
+Table 3: Relative lattice displacement at each catenation level. Amount of lattice displacement for catenated MOFs.
+catenation lattice1 lattice2 lattice3 lattice4
+cat0 (0,0,0) / / /
+cat1 (0,0,0) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) / /
+cat2 (0,0,0) ($\frac{1}{3}$,$\frac{1}{3}$,$\frac{1}{3}$) ($\frac{2}{3}$,$\frac{2}{3}$,$\frac{2}{3}$) /
+cat3 (0,0,0) ($\frac{1}{4}$,$\frac{1}{4}$,$\frac{1}{4}$) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) ($\frac{3}{4}$,$\frac{3}{4}$,$\frac{3}{4}$)
+Figure 6 shows the pairwise relationships of the CO2 capacities of hMOF structures at five pressures.
+We observe a strong correlation of CO2 capacities between 0.05 bar and 0.1 bar, with a Pearson’s correlation coefficient of 0.86.
+For other pairs of pressures, the CO2 capacities are only weakly correlated, with a decreasing correlation for larger pressure differences.
+Moreover, the distributions of CO2 capacities at all pressures exhibit long tails at high value ranges, which indicate that the majority of MOFs are low-performing and high-performing MOFs are uncommon.
+Figure 6: $\textrm{CO}_{2}$ capacities of hMOF structures at different adsorption pressures.
+Pairplot of $\textrm{CO}_{2}$ capacities of hMOF structures.
+The $(x,y)$ axes represent adsorption pressures.
+Different colors indicate MOFs with different node-topology pairs.
+Decomposing MOF linkers into molecular fragments
 The first component of the GHP-MOFassemble framework, Decompose, decomposes linkers from high-performing MOF structures in the hMOF dataset into their molecular fragments.
 As illustrated in Figure 7, this process applies the following three steps to a specified node-topology pair, which in the figure is the most frequent node-topology pair in the hMOF dataset, Zn tetramer-pcu.
-Note that for the pcu topology, one linker is orientated along each of the x, y, and z directions, and thus at most three linker types are possible.
+Note that for the pcu topology, one linker is orientated along each of the x, y, and z directions, and thus at most three unique linker types are possible.
   1. 1.
     Select: We select high-performing MOFs with the given node-topology pair from hMOF.
-We define here a high-performing MOF structure as one with CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of CO2 capacity.
+We define here a high-performing MOF structure as one with CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ at 0.1 bar, which corresponds to the top 5% of CO2 capacity of the hMOF dataset.
   2. 2.
-    Extract: We extract the linkers Simplified Molecular-Input Line-Entry System (SMILES) strings for each high-performing MOF identified in step (i), eliminate those for MOFs with more than three unique linker types, and assemble those that remain into a linker dataset.
+    Extract: We extract the linker Simplified Molecular-Input Line-Entry System (SMILES) strings for each high-performing MOF identified in Select step, eliminate those with more than three unique linker types, and assemble the remaining into a linker dataset.
   3. 3.
-    Fragment: We fragment the linkers produced in step (ii) to obtain their chemically relevant fragment-connection atom pairs, which we assemble into a molecular fragment dataset.
-The extraction of linkers in step (ii) is straightforward because the MOFid provided by hMOF for each MOF specifies the SMILES strings of its constituent inorganic and organic building blocks, including its linkers, as well as a format signature and a topology code [39].
+    Fragment: We fragment the linkers produced in Extract step to obtain their chemically relevant fragment-connection atom pairs, which we assemble into a molecular fragment dataset.
+The extraction of linkers in Extract step is straightforward because the MOFid of each hMOF structure specifies the SMILES strings of its constituent metal nodes, linkers, as well as a format signature and a topology code [52].
 Together, these elements uniquely define the topology and building blocks of a given MOF structure.
-Step (iii) use the Matched Molecular Pairs Algorithm (MMPA) [40] as implemented in DeLinker [41] to generate molecular fragments of a given molecule by breaking chemical bonds between atom pairs.
+Fragment step use MMPA [53] as implemented in DeLinker [54] to generate molecular fragments of a given molecule by breaking chemical bonds between atom pairs.
 We set the minimum number of connection atoms, the minimum fragment size, and the minimum path length to 3, 5, and 2, respectively.
@@
 The chemically relevant fragment-connection atoms pairs are then used to form the molecular fragment dataset.
-Figure 1: Schematic representation of Decompose, the first component of our GHP-MOFassemble framework.
-It consists of three steps, which we showcase for Zn tetramer-pcu MOFs.
-First, the Select step selects the Zn tetramer-pcu MOFs in hMOF that are high-performing.
-Then the Extract step extracts unique linkers from those MOFs.
+Figure 7: First component of the GHP-MOFassemble framework.
+Schematic representation of the Decompose component, which consists of three steps, which we showcase for Zn tetramer-pcu MOFs.
+First, the Select step selects the Zn tetramer-pcu MOFs in high-performing hMOF structures.
+Next, the Extract step extracts unique linkers from those MOFs.
 Finally, the Fragment step generates unique linker fragments.
 The color scheme of elements is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white.
-3.2 Generating new MOF structures
+See Table 1 for unique linker fragments of Zn tetramer node.
+Generating new MOF structures
 The Generate component employs the pre-trained diffusion model DiffLinker [31] to generate new MOF linkers, and then assembles those new linkers with one of three pre-selected nodes into new MOF structures in the pcu topology.
@@
 An example of these three steps for Zn tetramer-pcu MOFs is shown in Figure 8.
-3.2.1 Diffuse and Denoise.
-We apply the pre-trained diffusion model DiffLinker [14] to the molecular fragments output by the Decompose component to generate novel MOF linkers.
+Diffuse and Denoise
+We apply the pre-trained diffusion model DiffLinker [31] to generate new linkers based on the molecular fragments outputted by the Decompose component.
 This model connects the molecular fragments supplied as input (also known as context) with a specified number of sampled atoms.
-We vary the number of sampled atoms from five to 10, a range chosen because we found that with fewer than five sampled atoms, straight chains or branched chains may be obtained, while with more than 10, ring structures may be present.
-We found that the range of five to 10 sampled atoms resulted in novel and chemically diverse MOF linkers.
-DiffLinker applies a denoising process to determine the atomic species and Cartesian coordinates of sampled atoms by using a decoder neural network architecture named E(3)-Equivariant Graph Neural Network (EGNN[42].
+With less sampled atoms, straight chains or branched chains may be obtained, while with more sampled atoms, ring structures may be present.
+Moreover, during the denoising process, the species of the sampled atoms may change.
+In this work, we vary the number of sampled atoms from 5 to 10 to ensure that the generated linkers have a diverse set of chemistry and substructures.
+DiffLinker applies a denoising process to determine the atomic species and Cartesian coordinates of the sampled atoms by using a decoder neural network architecture named E(3)-Equivariant Graph Neural Network (EGNN) [55].
 This graph neural network predicts zero-mean centered Cartesian coordinates noise and one-hot encoding noise of atomic species, and accounts for equivariance due to translation, rotation, and reflection of molecules, i.e., Euclidean Group 3 (E(3)).
-Since the pre-trained DiffLinker model was trained on the GEOM dataset [43], it enables sampling of connection atoms with high chemical and structural diversity.
-DiffLinker outputs the 3D coordinates and the atomic species of heavy atoms in the linker molecules, rather than SMILES strings or 2D graphs.
-Since hydrogen atoms are implicit in the DiffLinker model, their 3D coordinates are not output by the model.
-To generate the spatial configurations of all-atom molecules, we employ openbabel to convert the DiffLinker generated molecules to SMILES strings, which in turn are used to generate the 3D configurations of linkers with explicit hydrogen atoms.
+In this work, we use the pre-trained DiffLinker model that was trained on the GEOM dataset [56], which contains 37 million molecular conformations for over 450,000 molecules.
+Such diverse training data enables DiffLinker to sample linker molecules with high chemical and structural diversity.
+Figure 8: Second component of the GHP-MOFassemble framework.
+The Generate component of the GHP-MOFassemble framework involves Diffuse and Denoise, Screen and Evaluate, and Assemble steps, as shown here for Zn tetramer-pcu MOFs (see Table 1 for unique linker fragments of Zn tetramer node).
+First, we generate new linkers via the Diffuse and Denoise step.
+We then add hydrogen atoms to ensure correct valency of the generated linkers.
+Next, we identify dummy atoms by replacing the two carbon atoms in the carboxyl groups of each generated linker with dummy atoms.
+Linkers with dummy atoms then undergo the Screen and Evaluate step, where we remove those with S, Br, and I elements, since these three elements are present in the GEOM dataset but not in the hMOF dataset.
+As a result, this step reduces the number of potential linkers to 5,162.
+The generated linkers’ molecular statistics are quantified using five metrics, including SAscore, SCscore, validity, uniqueness, and internal diversity.
+Finally, in the Assemble step, we build 40,000 new MOFs with Zn tetramer node in pcu topology.
+As before, the color scheme of elements is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white.
+See last column of Table 2 regarding this component.
+The outputs of DiffLinker are the 3D coordinates and the atomic species of heavy atoms in the linker molecules, rather than SMILES strings or 2D graphs.
+Since hydrogen atoms are implicit in the DiffLinker model, their 3D coordinates are not outputted by the model.
+To generate the spatial configurations of all-atom molecules, we employ openbabel to convert the DiffLinker generated molecules to SMILES strings, which in turn are used to generate 3D configurations of linkers with explicit hydrogen atoms.
 Openbabel uses distance-based heuristics to determine bond connectivity in a given molecule [57], which is critical to the assignment of the number of hydrogen atoms.
-After the hydrogen addition step, we identify the dummy atoms which contains information about how the linkers connect with metal nodes.
+Finally, after the hydrogen addition step, we identify the dummy atoms which contains information about how the linkers connect with metal nodes.
 The diffusion process involves two consecutive steps.
-The first adds Gaussian noise to the original data (i.e., x), yielding noisy input (i.e., $z_{t}$ at time t).
-Then, the denoising step applies the neural network-based noise removal operation.
+The first step adds Gaussian noise to the original data (i.e., x), yielding noisy input (i.e., $z_{t}$ at time t).
+Next, the denoising step applies neural network-based noise removal operation.
 Mathematically, the following Markovian properties and equations are satisfied:
@@
 Thus, invariant features such as one-hot and equivariant features such as atomic coordinates updates need to be considered [31].
-3.2.2 Screen and Evaluate.
-To ensure consistency of elements in hMOF linkers and generated linkers, we manuallyfilter out linkers that contain elements not present in the hMOF dataset.
-Since the pre-trained DiffLinker model we used in this work was trained on the GEOM dataset, a total of nine heavy elements may be sampled, including C, N, O, F, P, S, Cl, Br, and I.
-Among these elements, S, Br, and I do not appear in the hMOF dataset, therefore we remove linkers that contain these three elements.
+Screen and Evaluate
+To ensure consistency of elements in hMOF linkers and AI-generated MOF linkers, we manually filter out generated linkers that contain elements not present in the hMOF dataset.
+Since the pre-trained DiffLinker model we used in this work was trained on the GEOM dataset [56], a total of nine heavy elements may be sampled, including C, N, O, F, P, S, Cl, Br, and I.
+Among these elements, S, Br, and I do not appear in the hMOF dataset, therefore we remove generated linkers that contain these three elements.
 For each molecular fragment, we then perform sampling 20 times.
-Each sampling step may yield a different set of molecules because of the random nature of the denoising process.
-The probabilistic nature of the diffusion model allows it to generate linkers from an extensive linker design space beyond that of the hMOF dataset.
+Each time the model samples a different molecule may be obtained because of the random and probabilistic nature of denoising process.
+The probabilistic nature of diffusion model enables it to generate linkers from an extensive linker design space beyond that of the hMOF dataset.
 We then use five metrics to evaluate the quality of the remaining linkers.
@@
 For both SAscore and SCscore, the higher the values are, the more difficult it is to synthesize the linker, hence less desirable.
-To evaluate the capability of GHP-MOFassemble to generate a valid, novel, and diverse set of linkers, we leverage the MOSES framework [48] to compute three additional metrics: the fraction with valid SMILES strings (validity), the fraction that are unique (uniqueness), and their dissimilarity (internal diversity).
-We compute the fraction of linkers that are unique by comparing their canonical SMILES strings, and use their average pairwise molecular Tanimoto dissimilarity [49] as measures of diversity.
-We compute the diversity of the generated linker set by using two internal diversity scores $(\textrm{IntDiv}_{1}(G),\textrm{IntDiv}_{2}(G))$.
-We present the results of these metrics in Section 4.2.
-Figure 2: The Generate component of the GHP-MOFassemble framework involves Diffuse and Denoise, Screen, and Assemble steps, as shown here for Zn tetramer-pcu MOFs.
-First, we generate new linkers via the Diffuse and Denoise step.
-We then add hydrogen atoms to ensure correct valency of the generated linkers.
-Next, we identify dummy atoms by replacing the two carbon atoms in the carboxyl groups of each generated linker with dummy atoms.
-Linkers with dummy atoms then undergo the Screen step, which removes linkers with S, Br and I.
-These three elements are present in the GEOM dataset but not in hMOF dataset, and thus we filter out linkers with these three elements.
-As a result, this step reduces the number of potential linkers to 5,162.
-The generated linkers’ molecular statistics are quantified using five metrics, including SCscore, SAscore, validity, uniqueness, and internal diversity.
-Finally, in Assemble step, we build 40,000 new MOFs with the Zn tetramer node in pcu topology.
-As before, the color code used for the atoms shown is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white.
-3.2.3 Assemble.
-Once we have generated linkers that meet the five metrics mentioned above, we can assemble them by connecting the node and linker building blocks.
-To construct MOFs, we need to guide this assembly process.
-We do that by using dummy atoms, which indicate the points where the building blocks are to be connected.
+Specifically, the SAscore [60] is defined as the difference of fragment score and complexity penalty:
+$\mathrm{SAscore}=\mathrm{fragmentScore}-\mathrm{complexityPenalty},$
+where fragment score is calculated by averaging over the contributions of non-zero elements of the Morgan fingerprint for all molecular fragments.
+The complexity score is calculated based on five components:
+$\mathrm{complexityPenalty=-sizePenalty-stereoPenalty-spiroPenalty-bridgePenalty-macrocyclePenalty}.$
+The size penalty increases with the number of atoms:
+$\mathrm{sizePenalty=nAtoms^{1.005}-nAtoms},$
+The stereo penalty is calculated based on the number of chiral centers:
+$\mathrm{stereoPenalty=log(nChiralCenters+1)},$
+The spiro penalty is calculated based on the number of spiro centers, or atoms that connect two rings together:
+$\mathrm{spiroPenalty=\log(nSpiro+1)},$
+The bridge penalty term is calculated based on the number of bridgehead atoms, or atoms that belong to two or more rings:
+$\mathrm{bridgePenalty=\log(nBridgeheads+1)},$
+Lastly, if macrocycles are present, the macroPenalty is:
+[…]
+otherwise, the macroPenalty is zero.
+The SCscore [60] is calculated by using a neural network trained on 12 million reactions from the Reaxys database.
+The synthetic complexity of each molecule is rated from 1 to 5, with 1 being the easiest and 5 the most difficult to synthesize.
+To evaluate the capability of GHP-MOFassemble to generate a valid, novel, and diverse set of linkers, we leverage the MOSES framework [61] to compute three additional metrics: the fraction of valid linker SMILES strings (validity), the fraction of unique linker SMILES strings (uniqueness), and the dissimilarity of linkers (internal diversity).
+Validity measures whether atoms in generated molecules have the correct valency, whereas uniqueness quantifies the percentage of any molecule that is different from the rest of molecules.
+We show in Figure 9 the synthetic accessibility score (SAscore) and synthetic complexity score (SCscore) values for the remaining linkers.
+We observe that as the number of sampled atoms increases, both distributions generally shift to the right, with the exception of the SAscore distribution with six sampled atoms, which is to the left of that with five atoms.
+This general trend indicates that linkers become harder to synthesize as the molecules become bigger.
+This result is expected because as more atoms are sampled, more complex substructures may be present.
+We note that no linker has zero or very large SAscore or SCscore, values that would indicate unsynthesizability.
+Figure 9: Synthesizability of DiffLinker-generated MOF linkers with dummy atoms.
+a–c Distributions of synthetic accessibility score (SAscore); and d–f synthetic complexity score (SCscore) of DiffLinker-generated MOF linkers with dummy atoms.
+We indicate both node-topology pairs, and the number of sampled atoms, from 5 to 10 inclusive.
+We show in Table 4 the validity, uniqueness, and internal diversity metrics, grouped by the number of sampled atoms and node-topology pairs of the corresponding MOFs.
+The validity column confirms that all generated linkers are valid.
+The last three columns, when reviewed from top to bottom, reveal a considerable increase in linker uniqueness and a more modest increase in internal diversity as the number of sampled atoms increases.
+High uniqueness values indicate that our model is generating non-duplicate molecules, whereas high internal diversity values indicate that the generated molecules are chemically diverse.
+Internal diversity (which indicates how dissimilar a specific linker is to the rest of the population) is computed using two internal diversity scores: IntDiv1 and IntDiv2 [61], also shown in Table 4.
+The increase in these metrics as more atoms are sampled is expected because the degrees of freedom of atomic species and their spatial coordinates increase with molecular size.
+Table 4: Statistical properties of AI-generated linkers.
+Statistics of AI-generated linkers that correspond to different MOF types (first column), number of sampled atoms (${\cal{N}}$ in the second column), number of carboxyl linkers (${\cal{C}}$, fourth column), and number of heterocyclic linkers (${\cal{H}}$, fifth column).
+In total, there are 12,305 linkers (see Table 2’s Total column, Element filter row) with identified dummy atoms.
+[…]
+unique
+$\rm IntDiv_{1}$ $\rm IntDiv_{2}$
+Cu paddlewheel-pcu 5 1,240 774 466 1
+0.496
+0.692 0.669
+Zn paddlewheel-pcu 5 1,184 775 409 1
+0.505
+0.709 0.686
+Zn tetramer-pcu 5 1,666 1,666 0 1
+0.514
+0.681 0.659
+Cu paddlewheel-pcu 6 1,117 761 356 1
+0.545
+0.718 0.695
+Zn paddlewheel-pcu 6 992 681 311 1
+0.554
+0.718 0.697
+Zn tetramer-pcu 6 1,532 1,532 0 1
+0.512
+0.693 0.672
+Cu paddlewheel-pcu 7 499 383 116 1
+0.919
+0.751 0.734
+Zn paddlewheel-pcu 7 453 329 124 1
+0.933
+0.758 0.742
+Zn tetramer-pcu 7 709 709 0 1
+0.866
+0.744 0.729
+Cu paddlewheel-pcu 8 330 245 85 1
+0.947
+0.759 0.742
+Zn paddlewheel-pcu 8 338 247 91 1
+0.943
+0.763 0.745
+Zn tetramer-pcu 8 550 550 0 1
+0.905
+0.746 0.732
+Cu paddlewheel-pcu 9 281 191 90 1
+0.979
+0.756 0.739
+Zn paddlewheel-pcu 9 257 175 82 1
+0.989
+0.763 0.746
+Zn tetramer-pcu 9 389 389 0 1
+0.959
+0.742 0.727
+Cu paddlewheel-pcu 10 235 165 70 1
+0.988
+0.769 0.750
+Zn paddlewheel-pcu 10 217 148 69 1
+1.000
+0.768 0.749
+Zn tetramer-pcu 10 316 316 0 1
+0.994
+0.750 0.737
+The internal diversity scores were calculated based on the Tanimoto distances [62] among all pairs of molecules, which are obtained by calculating the normalized Jaccard score of Morgan fingerprint bit vector between all pairs of AI-generated linkers.
+This yields similarities between a specific linker and the rest of the linkers in the linker pool, which are averaged out.
+Therefore, we end up with a metric showing each linker’s similarity compared with the population of generated linkers.
+We used the MOSES [61] framework to compute the internal diversity scores IntDiv1 and IntDiv2 [61] by using the relations:
+$\textrm{IntDiv}_{1}(G)=1-\frac{1}{|G|^{2}}\sum_{m_{1},m_{2}\in G}T_{d}(m_{1},m_{2})\,,$ (7)
+$\textrm{IntDiv}_{2}(G)=1-\sqrt{\frac{1}{|G|^{2}}\sum_{m_{1},m_{2}\in G}T_{d}(m_{1},m_{2})^{2}}\,,$ (8)
+where $T_{d}$ is the Tanimoto distance, which relates to the Tanimoto similarity ($T_{s}$) by:
+$T_{d}(m_{1},m_{2})=1-T_{s}(m_{1},m_{2}),$ (9)
+where $G$ is the generated set of molecules, $|G|$ is the size of that set, and $(m_{1},m_{2})$ is a pair of molecules in the set.
+To measure the similarity between the AI-generated linkers and hMOF linkers in high-performing MOF structures, we show in Figure 10 the distribution of maximum Tanimoto similarity between the two linker sets.
+For each unique linker in the 364 AI-generated high-performing MOF structures we calculated its Tanimono similarity with all of the unique linkers in the hMOF structures.
+The maximum value of the Tanimoto similarities gives a quantitative measure of how different the generated linkers are from the hMOF linkers.
+Results in Figure 10 indicate that most linkers in the predicted high-performing MOF structures are vastly different from those in hMOF, i.e., our GHP-MOFassemble framework generates novel MOF structures with chemically unique linkers.
+Figure 10: Similarity between AI-generated and hMOF linkers in high-performing MOF structures.
+Distribution of maximum Tanimono similarity, which measures the uniqueness of AI-generated linkers as compared to hMOF linkers for high-performing MOFs.
+The peak around 0.3 to 0.4 indicates that most generated linkers are just 30–40% similar to those in hMOF, i.e., our AI framework generates novel linkers not present in hMOF structures.
+On the other hand, the trailing heavy right tail above 0.4 Tanimoto similarity indicates that we are also able to generate linkers that are structurally similar to those present in hMOF, showing that GHP-MOFassemble enables generation of a diverse set of novel linkers.
+MOF Assembly
+Once we have generated linkers, we can then assemble them with metal nodes.
+To construct MOFs, we need to guide the assembly process.
+We do this by using dummy atoms, which indicate the points at which the building blocks are to be connected.
 In practice, this is done as follows.
@@
 An illustration of dummy atom identification of linkers containing heterocyclic groups is shown in the right panel of Figure 11.
+Figure 11: Assembly of AI-generated linkers with metal nodes.
+a Identification of dummy atoms for linkers containing carboxyl groups.
+The dummy atoms are found by substituting the carbon atoms in the carboxyl groups.
+The remaining oxygen and hydrogen atoms in the carboxyl groups are removed.
+b Identification of dummy atoms for linkers containing heterocyclic rings.
+The dummy atoms are found at nitrogen-metal bond distance from the terminal nitrogen atoms along the vectors pointing from the opposing carbon atoms to nitrogen atoms.
 More than half of the hMOF structures are catenated MOFs, i.e., MOFs with interpenetrated lattices.
@@
 We denote the four catenation levels in hMOF, with increasing number of interpenetrated lattices, as cat0, cat1, cat2 and cat3.
-To generate MOFs with high structural diversity, we applied the site translation method as implemented in Pymatgen [50] to generate MOFs with different catenation levels.
+To generate MOFs with high structural diversity, we applied site translation method as implemented in Pymatgen [63] to generate MOFs with different catenation levels.
+Figure 12 presents the building block combinations of the final six MOF candidates that passed all screening processes.
+For linkers, the corresponding molecules (without dummy atoms) are shown for ease of visualization.
 For catenated MOF structures, we keep the spacing between interpenetrated lattices the same, so as to ensure equal pore sizes.
-3.3 Applying regression model to estimate MOF $\textrm{CO}_{2}$ capacity
-The Predict component of the GHP-MOFassemble framework uses a regression model to estimate the CO2 capacity of newly generated pcu MOF structures.
-We use for this purpose a modified version of the Crystal Graph Convolutional Neural Network (CGCNN) model, developed in our previous work [18], which uses an adjacency list to format node and edge embeddings, rather than the adjacency matrix format of the original CGCNN, a change that enhances model training speed, training stability, and prediction accuracy.
-We trained this modified CGCNN model on all hMOF structures, with a target property of CO2 capacity at 0.1 bar.
-We create our ensemble regression model by training the modified CGCNN model three times and averaging the results.
-We applied the resulting model to the assembled pcu MOF structures produced by the Generate step of Section 3.2 to identify those with a predicted CO2 capacity higher than 2 mmol/g.
-These MOFs (see Section 4.4.2) are candidates for further investigation by, for example, by performing CO2 capacity validation using Grand Canonical Monte Carlo (GCMC) simulations.
-4 Results
-By combining the generative power of diffusion models and the computing power of supercomputers, the GHP-MOFassemble framework enables rapid generation and assembly of chemically diverse MOF structures.
-Using the hMOF dataset as a starting point, our framework accelerates the computational design of high-performing MOF structures for carbon capture by high-throughput screening of novel MOF structures with high CO2 capacities.
-Our key findings are summarized below.
-4.1 Analysis of the hMOF dataset
-The three most frequent node-topology pairs in the hMOF dataset, accounting for around 74% of its MOFs, are the Cu paddlewheel-pcu, Zn paddlewheel-pcu, and Zn tetramer-pcu structures shown in the left panel of Figure 3.
-As summarized in Table 1, Zn tetramer-pcu is the most abundant, followed by Cu paddlewheel-pcu and then Zn paddlewheel-pcu.
-The cumulative distribution functions of the CO2 capacities of these three types of MOF structure are shown in the right panel of Figure 3.
-Only the 78,238 MOFs from hMOF with correctly parsed MOFids and valid SMILES were used for analysis.
-We gather the molecular fragments numbers in Table 1, produced through the Fragment step described in Section 3.1.
-In Table 1 high-performing MOFs with three parsed linkers were selected, and their unique linkers parsed by using the MMPA (Matched Molecular Pairs Algorithm) algorithm.
-In Table 1, the number of output molecular fragment conformers (last column) is much less than the number of unique linker SMILES (second to last column) for two reasons.
-First, around 56% of linkers did not pass the valency check, which may be due to the intrinsic limitations in the parsing of MOF linkers.
-Second, around 90% of linkers that passed the valency check share similar molecular fragments, and thus many duplicated molecular fragments exist.
-The successfully parsed molecular fragment conformers are subsequently used for linker generation.
-Figure 3: Depictions of the most frequent node-topology pairs in hMOF (left panel) and cumulative distribution functions of the 0.1 bar CO2 capacities (right panel).
-Table 1: Data on the hMOF with most frequent node-topology pairs. PW, TM, and HP-MOF stand for paddlewheel, tetramer, and high-performing MOF, respectively.
-node-topology total MOFs HP-MOFs HP-MOFs with unique linker unique molecular
-three linkers SMILES fragment conformers
-Cu PW-pcu 29,714 1,458 1,016 3,330 180
-Zn PW-pcu 28,529 1,314 834 3,221 162
-Zn TM-pcu 43,874 2,129 1,388 3,265 198
-Figure 4: Empirical cumulative distribution functions of the 0.1 bar CO2 capacities of hMOF structures at different catenation levels. The x axis is capped at 4 mmol/g to preserve details.
-Figure 4 shows the empirical cumulative distribution functions of the CO2 capacities of hMOF structures at different catenation levels.
-Therein, we observe that a higher percentage of catenated MOFs (cat1, cat2 and cat3) are high-performing, as compared to the uncatenated MOFs (cat0).
-This result confirms that catenation is an important factor when designing new MOF structures with high CO2 capacity.
-In Table 5, we summarize the amount of origin shift for each of the four levels of catenation.
-Figure 5 shows the pairwise relationships of the CO2 capacities of hMOF structures at the five pressures.
-We observe a strong correlation of CO2 capacities between 0.05 bar and 0.1 bar, with a Pearson’s correlation coefficient of 0.86.
-For other pairs of pressures, the CO2 capacities are only weakly correlated, with a decreasing correlation for larger pressure differences.
-Moreover, the distributions of CO2 capacities at all pressures exhibit long tails at high value ranges, which indicate that the majority of MOFs are low-performing and high-performing MOFs are uncommon.
-Figure 5: Pairplot of the CO2 capacities of hMOF structures at different adsorption pressures.
-The x- and y-axis are adsorption pressures.
-Different colors indicate MOFs with different node-topology pairs.
-4.2 Linker generation and evaluation
-For all three MOF types, we start with 540 molecular fragments extracted from high-performing hMOF structures, and use DiffLinker to generate a large pool of 64,800 linkers with the number of sampled connection atoms ranging from 5 to 10.
-For each linker, sampling is performed 20 times, as described in Section 3.2.
-Since these linkers only contain heavy atoms, to generate all-atom linker molecules, we apply openbabel to add hydrogen atoms, which results in 56,257 linkers.
-Next, dummy atom identification is performed to generate information that enables assembly with metal nodes.
-A total of 16,162 linkers with generated dummy atoms are generated.
-These linkers are passed through the element filter, which removes linkers that contain S, Br and I elements, further reducing number to 12,305 linkers.
-The details of the number of linkers in each step are summarized in Table 7.
-We show in Figure 6 the synthetic accessibility score (SAscore) and synthetic complexity score (SCscore) values, defined in Section 3.2.2, for the remaining linkers.
-We observe that as the number of sampled atoms increases, both distributions generally shift to the right, with the exception of the SAscore distribution with six sampled atoms, which is to the left of that with five atoms.
-The general trend indicates that linkers become harder to synthesize as the molecules become bigger.
-This result is expected because as more atoms are sampled, more complex substructures may be present.
-We note that no linker has zero or very large SAscore or SCscore, values that would indicate unsynthesizability.
-Figure 6: Distributions of synthetic accessibility score (SAscore, top row) and synthetic complexity score (SCscore, bottom row) of DiffLinker-generated MOF linkers with dummy atoms.
-The subfigure titles are node-topology pairs, and the colors listed in the legend indicate the number of sampled atoms, from five to 10 inclusive.
-We show in Table 2 the validity, uniqueness, and internal diversity metrics, defined in Section 3.2.2, grouped by number of sampled atoms and node-topology pairs of the corresponding MOFs.
-The validity column confirms that all generated linkers are valid.
-The last three columns, when reviewed from top to bottom, reveal a considerable increase in linker uniqueness and a more modest increase in internal diversity as the number of sampled atoms increases.
-High uniqueness values indicate that our model is generating non-duplicate molecules, whereas high internal diversity values indicate that the generated molecules are chemically diverse.
-Internal diversity (which indicates how similar/dissimilar a specific linker is to the rest of the population) is computed using two internal diversity scores: IntDiv1 and IntDiv2 [48], also shown in Table 2—see C for details on how these scores are computed.
-The increase in these metrics as more atoms are sampled is expected because the degrees of freedom of atomic species and their spatial coordinates increase with molecular size.
-Table 2: Statistics of the generated linkers that correspond to different MOF types (first column), number of sampled atoms (${\cal{N}}$ in the second column), number of carboxyl linkers (${\cal{C}}$, fourth column), and number of heterocyclic linkers (${\cal{H}}$, fifth column).
-In total, there are 12,305 linkers with identified dummy atoms.
+Figure 12: Optimal linkers and catenation levels for MOF assembly. Building blocks and catenation levels of the top six AI-generated MOF candidates.
+Screening and predicting properties of AI-generated assembled MOFs
+Here we describe the final Screen and Predict component of GHP-MOFassemble.
+After the MOF structures are assembled, a comprehensive screening workflow is developed to sift through the candidates with increasing levels of computational cost and confidence level:
+  • Inter-atomic Distance Check (screen)
+  • Pre-simulation Check (screen)
 […]
-unique
-$\rm IntDiv_{1}$ $\rm IntDiv_{2}$
-Cu paddlewheel-pcu 5 1,240 774 466 1
-0.708
-0.789 0.769
-Zn paddlewheel-pcu 5 1,184 775 409 1
-0.692
-0.783 0.762
-Zn tetramer-pcu 5 1,666 1,666 0 1
-0.734
-0.739 0.721
-Cu paddlewheel-pcu 6 1,117 761 356 1
-0.764
-0.819 0.796
-Zn paddlewheel-pcu 6 992 681 311 1
-0.761
-0.809 0.787
-Zn tetramer-pcu 6 1,532 1,532 0 1
-0.766
-0.749 0.732
-Cu paddlewheel-pcu 7 499 383 116 1
-0.958
-0.836 0.820
-Zn paddlewheel-pcu 7 453 329 124 1
-0.929
-0.829 0.812
-Zn tetramer-pcu 7 709 709 0 1
-0.958
-0.782 0.770
-Cu paddlewheel-pcu 8 330 245 85 1
-0.992
-0.835 0.822
-Zn paddlewheel-pcu 8 338 247 91 1
-0.978
-0.831 0.817
-Zn tetramer-pcu 8 550 550 0 1
-0.983
-0.791 0.780
-Cu paddlewheel-pcu 9 281 191 90 1
-0.99
-0.837 0.824
-Zn paddlewheel-pcu 9 257 175 82 1
-0.993
-0.834 0.821
-Zn tetramer-pcu 9 389 389 0 1
-0.986
-0.794 0.784
-Cu paddlewheel-pcu 10 235 165 70 1
-0.996
-0.842 0.830
-Zn paddlewheel-pcu 10 217 148 69 1
-0.996
-0.837 0.826
-Zn tetramer-pcu 10 316 316 0 1
-0.995
-0.798 0.789
-4.3 MOF assembly
-We generate new MOF structures by assembling three randomly selected DiffLinker-generated linkers with one of the three most frequent nodes in the hMOF dataset.
-Random sampling of linkers ensures that the selected linkers cover a large design space.
-For each node-linker-topology combination, we considered four levels of catenation (cat0, cat1, cat2, cat3).
-We generated a total of 120,000 MOFs as follows: we have four catenation levels and three node candidates.
-The random sampling of 10,000 linkers for each catenation level-node candidate pair generates 120,000 total MOFs of different catenation-node-linker combinations.
-We provide more details in Section 3.2.3 on how these new MOFs were assembled.
-4.4 Regression model for MOF $\textrm{CO}_{2}$ capacity prediction
-4.4.1 Pre-training on hMOF structures
-We employ the modified version of the CGCNN model [18] as a predictive model of the CO2 capacity of MOF structures.
-We trained this model on 90% of the MOF structures in hMOF, along with their CO2 capacities at 0.1 bar, leaving 10% as a holdout set for testing.
-We trained the modified version of the CGCNN model three times to create an ensemble of models in order to mitigate biases in individual models.
-We summarize in Table 6 and Figure 12 the prediction errors of each model on the 10% test set and the distribution of standard deviations, respectively.
-Looking next at the ensemble model, we show in the left panel of Figure 7 its predictions for the 10% test set.
-The ensemble model has a mean absolute error (MAE) of 0.093 mmol/g on the test set.
-Using CO2 capacity of 2 mmol/g as a threshold, we repurposed our predictive model as a classifier to categorize MOFs into low and high performers (with predicted CO2 capacities below and above the threshold, respectively) and high performers.
+  • Structure Validation of the Assembled MOFs (screen)
+  • Grand Canonical Monte Carlo Simulation (screen and predict)
+Inter-atomic Distance Check
+A pairwise distance matrix $M_{n\times n}$ is extracted from a MOF’s CIF file. The matrix element $M_{i,j}$ denoting the Euclidean distance between atom $A_{i}$ and atom $A_{j}$ is:
+$M_{i,j}=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}+(z_{i}-z_{j})^{2}}\,,$ (10)
+where $1\leq i\leq n$ and $1\leq j\leq n$, and $(x_{i},y_{i},z_{i})$ and $(x_{j},y_{j},z_{j})$ are the Cartesian coordinates of atoms $A_{i}$ and $A_{j}$.
+For an arbitrary pair of atoms $A_{i}$ and $A_{j}$, their chemical symbols can be denoted as $E_{i}$ and $E_{j}$, respectively.
+The bond length between atoms $A_{i}$ and $A_{j}$ is $\delta(E_{i},E_{j})$.
+The minimum allowed inter-atomic distance between the these two atoms can be estimated by calculating the infimum of all experimental bond lengths between these two element types, i.e.: $\inf{(\delta(E_{i},E_{j}))}$, which can be queried from the OChemDb [64] database.
+All values of $M_{i,j}\ (i\neq j)$ are compared against $\inf{(\delta(E_{i},E_{j}))}$, and any occurrence of $M_{i,j}<\inf{(\delta(E_{i},E_{j}))}\ (i\neq j)$ will result in the MOF candidate being discarded.
+Entries with $i=j$ (i.e., the diagonal entries of the distance matrix $M_{i,i}$) are not examined as they should always be zero, because the distance between an atom and itself is zero.
+We used Python 3.10’s built-in multiprocessing library to help launch the inter-atomic distance check on different MOF structures in parallel.
+Pre-simulation Check
+For each atom in the MOF structure, the pre-simulation check workflow can be described as the following:
+  1. 1.
+    The atom’s element should be recognized as one of the supported elements in the UFF4MOF parameter set,
+  2. 2.
+    The atom’s neighboring atoms that are within the allowed bonding distance are examined for chemistry validity.
+We conducted this process with the help of the cif2lammps library. Python’s multiprocessing library is also used to speed up this process over many structures in parallel.
+Predicting $\textrm{CO}_{2}$ capacity of MOF via pre-trained regression model
+The GHP-MOFassemble framework uses an ensemble of regression models to estimate the CO2 capacity of newly generated pcu MOF structures.
+We use for this purpose a modified version of the CGCNN model, developed in our previous work [37], which adopts an adjacency list to format node and edge embeddings, rather than the adjacency matrix format of the original CGCNN, a change that enhances model training speed, training stability, and prediction accuracy.
+To fine-tune this AI model, we used MOF structures in the hMOF dataset, along with their CO2 capacities at 0.1 bar as input.
+We split the hMOF dataset into 80% training set, 10% validation set, and 10% test test.
+We independently trained three modified versions of the CGCNN model for 5,000 epochs with a batch size of 160, using Adam as the optimizer, at a learning rate of $10^{-4}$, and a weight decay rate of $2\times 10^{-5}$.
+Table 5 shows the R2 score, mean absolute error (MAE), and root mean squared error (RMSE) of the three AI models ensemble on the 10% test set.
+Figure 13 shows the distribution of the standard deviations of ensemble model predictions.
+We treat the standard deviation of predictions from the three models as a measure of model uncertainty, which we view as arising from the model’s difficulty in learning certain data points.
+This uncertainty is also known as epistemic uncertainty.
+By employing an ensemble of models and averaging their prediction results, we can increase prediction accuracy, which is thoroughly and extensively explored in the machine learning community [65, 66].
+We chose the threshold for the standard deviation filter to be 0.2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ because it is sufficiently small (with 96% of data points below the threshold) that for low CO2 capacity predictions, the predicted values across the ensemble are similar.
+On the other hand, for high CO2 capacity predictions, it also ensures that the outliers (i.e., predicted values across the ensemble are vastly different) are filtered out, therefore minimizing the overall error.
+In Table 6 we also notice that most of the high-performing MOFs are cat2 and cat3.
+This is consistent with our preliminary analysis on catenation levels and distributions of CO2 adsorption capacity in Figure 5.
+Table 5: Statistical properties of AI ensemble to characterize MOF’s performance. Statistics of AI ensemble, including R2 score, mean absolute error (MAE), and root mean squared error (RMSE).
+Model R2 MAE RMSE
+Model1 0.932 0.098 0.171
+Model2 0.937 0.100 0.170
+Model3 0.936 0.099 0.170
+Figure 13: Standard deviation of AI model ensemble to estimate MOFs’ CO2 capacity.
+Distribution of the standard deviation of AI ensemble predictions.
+Predictions with standard deviation of less than 0.2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ consist of around 96% of the test set.
+This shows that our three independently trained AI models are in great agreement.
+The remaining 4% with larger than 0.2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ standard deviation implies that these are the data points which may be difficult to predict due to errors in target property or extra information other than atomic species and periodic neighbor being necessary for accurate predictions.
+Table 6: Catenation level of high-performing MOFs. Numbers of predicted high-performing MOFs found in the hMOF dataset.
+node-topology cat0 cat1 cat2 cat3
+Cu paddlewheel-pcu 0 1 38 26
+Zn paddlewheel-pcu 0 0 66 23
+Zn tetramer-pcu 0 0 53 157
+We show in the left panel of Figure 14 the scatter plot of the ensemble model predictions for the 10% test set (from hMOF), which has a MAE of 0.093 m $\textrm{m\,mol}\,\textrm{g}^{-1}$.
+Using CO2 capacity of 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ as a threshold, we repurposed our predictive model as a classifier to categorize MOFs into low and high performers, with predicted CO2 capacities below and above the threshold, respectively.
 Using this scheme, the confusion matrix for identifying low and high performers is shown in the right panel of Figure 14.
-We conclude from the confusion matrix that the pre-trained model classifies both low and high performers with high accuracy, with 98.4% (13551 out of 13765) of test samples correctly classified.
-As the test set is heavily imbalanced, with many more low performers than high performers, we also calculated the balanced accuracy of classification [51] (the sum of true positive rate and true negative rate divided by 2), obtaining a value of 90.7%.
-The majority of the 214 misclassified samples lie close to the decision boundaries, as shown in the red and purple regions of the left panel of Figure 7.
-Therefore, we also conclude that the ensemble model is capable of differentiating low and high performers.
-Figure 7: Left panel, predictive performance of the ensemble model on the 10% test set.
-Right panel, confusion matrix of the ensemble model in classifying low and high performers in the test set.
-The classification threshold is 2 mmol/g.
-4.4.2 Predictions for newly generated MOF structures
-When using the ensemble model to infer the CO2 capacities of newly generated MOF structures, we employ the average of the predictions made by the three independent models as the predicted capacity.
-In addition, we filter out structures for which the standard deviation of the predicted values obtained from the three models is larger than 0.2 mmol/g.
-We treat the standard deviation of predictions from the three models as a measure of the uncertainty of model predictions, which we view as arising from the model’s difficulty in learning certain data points.
-We chose the threshold for the standard deviation filter of 0.2 mmol/g because it is sufficiently small (with 96% of data points below the threshold) that for low CO2 capacity predictions, the predicted values across the ensemble are similar.
-On the other hand, for high CO2 capacity predictions, it also ensures that outlier values (e.g., predictions made by the three models are vastly different) are filtered out, therefore minimizing the overall error.
-We show in Figure 8 the empirical cumulative distribution functions of both the hMOF (dashed lines) and predicted (sold lines) CO2 capacities.
-We observe that there are more high-performing cat2 and cat3 MOFs among the generated structures than in the hMOF dataset, as indicated by lower cumulative density values at 2 mmol/g (black dashed line).
-On the other hand, there are fewer high-performing cat0 and cat1 MOFs, as indicated by higher cumulative density values at 2 mmol/g.
-This result indicates that within the generated MOF design space, there are more high-performing candidates with higher catenation levels compared to lower catenation levels.
-Figure 8: Comparison of empirical cumulative distribution functions of the predicted CO2 capacities of generated and hMOF MOFs at different catenation levels.
-Out of the 120,000 hypothetical MOF structures (see Section 4.3 for MOF assembly details), a total of 6,020 were predicted to be high-performing MOFs with CO2 capacity higher than 2 mmol/g at 0.1 bar.
-Categorization of the predicted high-performing MOFs by node-topology pair and catenation level (see Table 3) shows that the majority of Cu paddlewheel-pcu and Zn tetramer-pcu are cat2 and cat3, whereas the majority of high-performing Zn paddlewheel-pcu are cat1 and cat2.
-Table 4 shows the building blocks and catenation levels in the top five pcu MOF candidates.
-For linkers, the corresponding molecules (without dummy atoms) are shown for ease of visualization.
-Note that some substructures appear more often than others.
-For example, ring structures appear in most of the linkers in the top five candidates, and many of the DiffLinker-generated molecular substructures (in between the terminal fragments) also contain ring structures.
-The frequent occurrence of rings in linkers of high performing MOFs may be due to the presence of aromatic linkers, as reported by phenylene et al.
-[52].
-We compared the chemical similarity of the linkers in predicted high-performing MOF structures with those in high-performing hMOF structures.
-As shown in Figure 13, we find that most linkers in the predicted high-performing MOF structures are vastly different from those in hMOF, indicating that the GHP-MOFassemble framework generate novel MOF structures with chemically unique linkers.
-Table 3: Numbers of predicted high-performing MOFs found in the hMOF dataset.
-node-topology cat0 cat1 cat2 cat3
-Cu paddlewheel-pcu 0 170 1,184 628
-Zn paddlewheel-pcu 9 789 929 395
-Zn tetramer-pcu 0 3 794 1,119
-Table 4: Details of the five pcu MOF candidates generated by GHP-MOFassemble with the highest predicted CO2 capacities.
-Their stability was confirmed using molecular dynamics simulations, and their 3D structures are shown in Figure 10.
-5 Validation of AI-generated MOFs with molecular dynamics simulations
-We now examine the stability and porous properties of the 6,020 MOFs described in the previous section by using molecular dynamics (MD) simulations.
-We use the LAMMPS [53] code, version 28 Mar 2023 Update 1, distributed from the conda-forge channel, to perform MD simulations that relax the AI-generated MOF structures.
-An adapted version of cif2LAMMPS [54] is used to generate LAMMPS input files automatically.
-For each MOF, a 2×2×2 supercell structure is generated, and an NPT simulation is performed, i.e., we use an isothermal-isobaric ensemble with a constant particle number $N$, a pressure $p$ fluctuating around an equilibrium value, $\langle p\rangle=1\textrm{atm}$, and a temperature $T$ fluctuating around an equilibrium value $\langle T\rangle=300\textrm{K}$.
-The cell lengths and angles of the triclinic simulation box are allowed to relax.
-We run the NPT simulations for 200,000 steps with a timestep of 1 femtosecond.
-The relative changes of structure density before and after the simulation are inspected to evaluate how well the porous characteristics are maintained throughout the simulation.
-This process produces 102 MOFs whose structures stayed intact during the MD simulation.
-Next, we use a $<$1% change in density as the threshold to determine whether a MOF structure is stable or not.
-Among the 102 MOF candidates, 18 have change in density below the 1% threshold.
-Figure 9 presents the top five AI-generated, stable MOF structures that passed the threshold for structure determination—see Table 4 for combinations of their building blocks and Figure 10 for 3D visualization of their crystal structures.
-Figure 9: Top five AI-generated MOFs with structure density changed less than 1% during molecular dynamics simulations.
-Their CO2 capacities are higher than 96.9% of hMOF structures in the hMOF database.
-Figure 10: 3D visualization of the crystal structures of top five AI-generated, stable MOF candidates that retained their porous properties throughout molecular dynamics simulations.
-Atoms are presented using the color code: carbon in grey, nitrogen in dark blue, fluorine in cyan, zinc in purple, hydrogen in white, and lithium in green.
-6 Conclusion
-We have introduced GHP-MOFassemble, an AI-driven discovery framework for high-throughput generation of novel high-performing MOF structures for carbon capture.
-The GHP-MOFassemble framework employs a diffusion model, DiffLinker, to generate new, unique, and chemically diverse MOF linkers, which it then assembles with one of three pre-selected metal nodes into MOFs in the pcu topology.
-Out of 120,000 MOF structures generated by GHP-MOFassemble, 6,020 were predicted to be high-performing MOFs with CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of the hMOF dataset.
-We relaxed each of these 6,020 MOFs in a 200-picosecond NPT molecular dynamics simulation with the UFF4MOF force field [55, 56], and found that 102 structures did not collapse after molecular dynamics simulations, among which 18 were identified to be stable structures with less than 1% change in density.
-We have deployed and extensively tested GHP-MOFassemble on computers at the Argonne Leadership Computing Facility and at the National Center for Supercomputing Applications, with the intent of providing scalable and computationally efficient AI tools to accelerate the modeling and discovery of novel MOF structures.
+We conclude from the confusion matrix that the pre-trained model classifies both low and high performers with high accuracy, with 98.4% (13,551 out of 13,765) of test samples (of hMOF) correctly classified.
+As the test set is heavily imbalanced, with many more low performers than high performers, we also calculated the balanced accuracy of classification [67] (the sum of true positive rate and true negative rate divided by two for binary classification or average of recalls for multi-class classification), obtaining a value of 90.7%.
+Since the majority of the 214 misclassified samples lie close to the decision boundaries, as shown in the red and purple regions of the left panel of Figure 14, we conclude that the ensemble model is capable of differentiating low and high performers.
+Figure 14: Performance of AI ensemble to measure CO2 capacity.
+a Predictive performance of the ensemble model on the 10% test set of hMOF.
+b Confusion matrix of the ensemble model in classifying low and high performers in the test set.
+MOFs with real/predicted CO2 capacity lower than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ are identified as low performers, and MOFs with real/predicted CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ are identified as high performers.
+We show in Figure 15 the empirical cumulative distribution functions of the CO2 capacities of hMOF (dashed lines) structures and AI-generated MOF structures (solid lines) at different catenation levels.
+We observe that there are more high-performing cat2 and cat3 MOFs among the generated structures compared to the hMOF dataset, as indicated by lower cumulative density values at 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ (black dashed line).
+On the other hand, there are fewer high-performing cat0 and cat1 MOFs, as shown by higher cumulative density values at 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$.
+This result demonstrates that within the AI-generated MOF design space, there are more high-performing candidates with higher catenation levels compared to lower catenation levels.
+Figure 15: MOFs’ CO2 capacities in terms of catenation levels.
+Comparison of empirical cumulative distribution functions of the predicted CO2 capacities of generated and hMOF structures for a cat0, b cat 1, c cat2, and d cat 3.
+In summary, we use an ensemble of 3 AI models to identify pcu AI-generated MOFs produced by the Generate step (in addition to inter-atomic distance and pre-simulation checks) that have predicted CO2 capacities higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$.
+We use the ensemble mean plus standard deviation as the final prediction.
+These MOFs are then selected for further investigation with MD simulations.
+Structure Validation of the Assembled MOFs
+Pre-screened MOFs are checked for stability using MD simulations with LAMMPS [42], version stable release 2 August 2023 compiled with gcc 11.2.0 and OpenMPI 4.1.2.
+For each MOF, a 2x2x2 supercell of triclinic periodic boundary condition is created.
+A triclinic isothermal-isobaric ensemble (i.e., NPT) at 300 K and 1 atm is applied to the supercell structure.
+All lattice parameters: $a$, $b$, $c$, $\alpha$, $\beta$, $\gamma$ are allowed to equilibrate freely.
+A relatively short simulation with 400,000 steps is conducted for each MOF to allow for structural equilibrium with a step size of 0.5 femtoseconds.
+The potential energy $E_{\textrm{pot}}$ of the MOF structure can be expressed as:
+$E_{\textrm{pot}}=E_{\textrm{bond}}+E_{\textrm{angle}}+E_{\textrm{dihedral}}+E_{\textrm{improper}}+E_{\textrm{Vdw}}+E_{\textrm{Coul}},$ (11)
+where $E_{\textrm{bond}}$ is the bond stretch energy, $E_{\textrm{angle}}$ is the angle between bonds energy, $E_{\textrm{dihedral}}$ is the dihedral angle energy, $E_{\textrm{improper}}$ is the improper dihedral angle energy, $E_{\textrm{vdw}}$ is the nonbonded interaction energy, and $E_{\textrm{coul}}$ is the Coulombic interaction energy.
+Coupry et al.
+(2016) [40] stated that 76.5% of the CoRE structures [14] lattice parameters are within 5% change when they are simulated with the UFF4MOF parameters.
+A similar criterion is used here such that if any of MOF’s lattice parameters of triclinic cell changes are more than 5%, the structure is discarded in the screening process.
+This screening process provides a new level of confidence in terms of structural stability of the AI-generated MOFs.
+Figure 16 demonstrates the relative error of all lattice parameters of the top six MOF candidates during the MD simulation.
+Figure 16: Lattice parameter relative errors of the top 6 AI-generated MOFs.
+The top six AI-generated MOF candidates have change less than 5% in all lattice parameter during MD simulation.
+The first three lattice parameters are cell axes length and the latter three lattice parameters are cell plane angles.
+Their CO2 capacities are higher than 96.9% of MOF structures in the hMOF database.
+Grand Canonical Monte Carlo (GCMC) Simulations
+The void fraction of each candidate MOF is calculated using the RASPA [43]’s helium void fraction function.
+It helps us understand the porosity contribution toward MOF CO2 adsorption capacity.
+Next, PACMOF [68] is used to assign partial charges to MOFs.
+Using the pre-trained partial atomic charge model on the density derived electrostatic and chemical (DDEC) charge [69] data, the partial charges are assigned to atoms in MOF structures.
+After determining the partial charges, the MOF structures are assigned with UFF4MOF force field parameters.
+The GCMC simulations are conducted at 0.1 bar and 300 K with a step size of 0.5 fs/step.
+The MOF structures are under the rigid assumption, therefore only Van der Waals interactions and Coulombic interactions (i.e., non-bonded) are considered:
+$E_{\textrm{pot}}=E_{\textrm{Vdw}}+E_{\textrm{Coul}},$ (12)
+which can be expressed as [70]
+$E_{\textrm{pot}}=\sum_{i,j}{4\epsilon_{ij}\bigg{[}\Big{(}\frac{\sigma_{ij}}{r_{ij}}\Big{)}^{12}+\Big{(}\frac{\sigma_{ij}}{r_{ij}}\Big{)}^{6}\bigg{]}}+\frac{q_{i}q_{j}}{4\pi\epsilon_{0}r_{ij}},$ (13)
+where $\epsilon_{ij}$, $\sigma_{ij}$, and $r_{ij}$ are the well depth, zero position, and inter-atomic distance of a Lennard-Jones 12-6 interaction between atom $i$ and atom $j$; $q_{i}$ and $q_{j}$ are the electrostatic charges of atom $i$ and atom $j$; $\epsilon_{0}$ is the vacuum permittivity.
+Computational resources and performance
+We have deployed and extensively tested GHP-MOFassemble on computers at the ALCF and at the National Center for Supercomputing Applications (NCSA), with the intent of providing scalable and computationally efficient AI tools to accelerate the modeling and discovery of novel MOF structures.
 The tools introduced in this work may be readily fine-tuned and adapted to other available datasets beyond hMOF to enable accelerated design and discovery of novel MOF structures for carbon capture at industrial scale.
-7 Acknowledgments
+The training and inference of the generative model and the training of the regression model was conducted on 8-way NVIDIA A100 GPUs with FP32 mode.
+The inference of the regression model is conducted on a single NVIDIA A40 GPU with FP32 mode.
+CPU-based screening processes, including distance check, pre-simulation check, MD simulations, and GCMC simulations, are run on two-way AMD Epyc 7763 CPUs.
+We now present a breakdown of the average computational cost of each step of the GHP-MOFassemble workflow per MOF:
+  • MOF assembly: 0.46 seconds per core. We scaled up this analysis by multiprocessing on 28 cores.
+  • Distance check to ensure structural validity: 2.56 seconds per core. We scaled up this analysis by multiprocessing on 128 cores.
+  • Pre-simulation check to ensure chemical consistency: 19.98 seconds per core. We scaled up this analysis by multiprocessing on 128 cores.
+  • AI ensemble inference of CO2 adsorption capacity: 20.46 seconds. We used one NVIDIA A40 GPU.
+  • MD simulations to validate stability and chemical consistency: 658.40 seconds, or about $\sim 10$ minutes.
+We scaled up this analysis using between six to 14 MPI processes based on the number of atoms in the MOF.
+  • Detailed GCMC simulations: 21214.04 seconds, or about $\sim 6$ hours. This averaged number is based on simulations on one CPU core.
+A schematic representation of this benchmark analysis is presented in  Figure 17.
+Figure 17: Process time cost breakdown of GHP-MOFassemble.
+Benchmark analysis of the computational cost for each process of the GHP-MOFassemble framework.
+Notice that the AI component is hyperefficient.
+Data Availability
+The datasets generated during and/or analysed during the current study are available in the GitHub repository, https://github.com/hyunp2/ghp_mof/tree/main/utils.
+We also used the open source hMOF dataset [71], and the GEOM dataset [56].
+Acknowledgements (not compulsory)
 This work was supported by Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, provided by the Director, Office of Science, of the U.S.
@@
 Code availability
-The code and data are available upon reasonable request from the authors.
-ORCID IDs
-Hyun Park $0$-0001-5550-5610
-Xiaoli Yan $0$-0002-6512-2338
-Ruijie Zhu $0$-0001-9316-7245
-Eliu Huerta $0$-0002-9682-3604
-Santanu Chaudhuri $0$-0002-4328-2947
-Ian Foster $0$-0003-2129-5269
-Emad Tajkhorshid $0$-0001-8434-1010
-Appendix A Dummy atom identification
-Figure 11 illustrates the workflow used to identify the dummy atoms and to remove redundant atoms for linkers containing carboxyl groups (left panel) or heterocyclic rings (right panel).
-Figure 11: Left panel, identification of dummy atoms for linkers containing carboxyl groups.
-The dummy atoms are found by substituting the carbon atoms in the carboxyl groups.
-The remaining oxygen and hydrogen atoms in the carboxyl groups are removed.
-Right panel, identification of dummy atoms for linkers containing heterocyclic rings.
-The dummy atoms are found at nitrogen-metal bond distance from the terminal nitrogen atoms along the vectors pointing from the opposing carbon atoms to nitrogen atoms.
-Appendix B MOF catenation via site translation
-Catenated MOFs were generated by using the site translation method, which is achieved by displacing the reference lattice along the diagonal line of the unit cell.
-For all four catenation levels, the amount of relative lattice displacement is given in Table 5.
-The numbers are fractional displacements relative to the unit cell diagonal.
-Table 5: Amount of lattice displacement for catenated MOFs
-catenation lattice1 lattice2 lattice3 lattice4
-cat0 (0,0,0) / / /
-cat1 (0,0,0) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) / /
-cat2 (0,0,0) ($\frac{1}{3}$,$\frac{1}{3}$,$\frac{1}{3}$) ($\frac{2}{3}$,$\frac{2}{3}$,$\frac{2}{3}$) /
-cat3 (0,0,0) ($\frac{1}{4}$,$\frac{1}{4}$,$\frac{1}{4}$) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) ($\frac{3}{4}$,$\frac{3}{4}$,$\frac{3}{4}$)
-Appendix C Linker evaluation
-Recall from Section 4.2 that we use the synthetic accessible score (SAscore) and synthetic complexity score (SCscore) metrics to confirm that all linkers are synthesizable.
-Table 2 provides additional data on the generated linkers.
-We also provide results for the validity and uniqueness metrics, and the internal diversity scores IntDiv1 and IntDiv2 [48].
-The internal diversity scores were calculated based on the Tanimoto distances among all pairs of molecules, which are obtained by calculating the normalized Jaccard score of Morgan fingerprint bit vector between all pairs of generated linkers.
-This yields similarities between a specific linker and the rest of the linkers in the linker pool, which are averaged out.
-Therefore, we end up with a metric showing each linker’s similarity compared with the population of generated linkers.
-We used the MOSES [48] framework to compute these two scores using the relations:
-$\textrm{IntDiv}_{1}(G)=1-\frac{1}{|G|^{2}}\sum_{m_{1},m_{2}\in G}T_{d}(m_{1},m_{2})\,,$ (7)
-$\textrm{IntDiv}_{2}(G)=1-\sqrt{\frac{1}{|G|^{2}}\sum_{m_{1},m_{2}\in G}T_{d}(m_{1},m_{2})^{2}}\,,$ (8)
-where $G$ is the generated set of molecules, $|G|$ is the size of molecule set, $(m_{1},m_{2})$ represents a pair of molecules in the molecule set.
-$T_{d}$ is the Tanimoto distance, which relates to the Tanimoto similarity ($T_{s}$) by:
-$T_{d}(m_{1},m_{2})=1-T_{s}(m_{1},m_{2})$ (9)
-Appendix D Statistics and distribution of the standard deviation of the ensemble of pre-trained models
-Table 6 shows the R2 score, mean absolute error (MAE), and root mean squared error (RMSE) of the three pre-trained models used for creating the ensemble model.
-Figure 12 shows the distribution of the standard deviations of ensemble model predictions.The threshold 0.2 mmol/g serves as a criterion for assessing whether the predictions made by the three pre-trained models are in agreement.
-The ensemble model predictions with standard deviation above the threshold were discarded.
-Table 6: Statistics of the three pre-trained models
-Model R2 MAE RMSE
-Model1 0.932 0.098 0.171
-Model2 0.937 0.100 0.170
-Model3 0.936 0.099 0.170
-Figure 12: Distribution of the standard deviation of ensemble model predictions.
-The ensemble model predictions with standard deviation of less than 0.2 mmol/g consist of around 96% of the test set.
-This shows that our three independently trained models are in great agreement for most of the unobserved test set, showing robustness of our model.
-The remaining 4% with larger than 0.2 mmol/g standard deviation implies that these are the data points which may be difficult to predict due to errors in target property or extra information other than atomic species and periodic neighbor is necessary for accurate prediction.
-Appendix E Linker similarity for high-performing MOF structures
-To measure the similarity between the generated linkers and hMOF linkers in high-performing MOF structures, we show in Figure 13 the distribution of maximum Tanimoto similarity between the two linker sets.
-For each unique linker in the predicted high-performing MOF structures generated by the GHP-MOFassemble framework, we calculate its Tanimono similarity with all of the unique linkers in the hMOF structures.
-The maximum value of the Tanimoto similarities gives a quantitative measure of how different the generated linkers are from the hMOF linkers.
-Figure 13: Distribution of maximum Tanimono similarity between generated and hMOF linkers in high-performing MOF structures.
-The peak around 0.3 to 0.4 indicates that most generated linkers are just 30–40% similar to those in hMOF–showing that we are generating novel linkers not present in hMOF structures.
-On the other hand, the trailing heavy right tail above 0.4 Tanimoto similarity indicates that we are also able to generate linkers that are structurally similar to those present in hMOF, showing that GHP-MOFassemble enables generation of a diverse set of novel linkers.
-Appendix F Regression model training details
-We independently trained three modified versions of the CGCNN model for 5,000 epochs with a batch size of 160, and with optimizer, learning rate, and weight decay of torch_adam, 1e-4, and 2e-5, respectively.
-Appendix G Linker statistics
-Table 7: Statistical summary of the number of linkers after each step categorized by the corresponding MOF types.
-PW and TM stand for paddlewheel and tetramer, respectively.
-Element filter means that linkers with S, Br and I elements are removed.
-Method Total Cu PW-pcu Zn PW-pcu Zn TM-pcu
-DiffLinker 64,800 21,600 19,440 23,760
-Hydrogen addition 56,257 18,979 17,126 20,152
-Dummy atom identification 16,162 4,964 4,450 6,748
-Element filter 12,305 3,702 3,441 5,162
+The scientific software and data used in this article are readily available in GitHub at https://github.com/hyunp2/ghp_mof/tree/main.
+Author contributions statement
+E.A.H.
+envisioned and led this work, and guided the connection between generative AI with MD and GCMC simulations to create novel, chemically consistent MOFs at scale in high performance computing platforms.
+H.P.
+adapted techniques from drug design into MOF discovery with generative AI, producing a pool of candidates of MOF components, and also trained and inferred assembled MOFs’ CO2 adsorption capabilities using a graph neural network model.
+X.Y.
+developed codes to decompose MOF into linkers and nodes, as well as to assemble predicted linkers and nodes to make hypothetical MOFs, and also conducted large-scale MD simulations, with LAMMPS and GCMC to identify stable and synthesizable AI-predicted MOFs.
+R.Z.
+analyzed hMOF database to understand MOF composition statistics, tested combined generative AI and graph models to predict novel AI-generated MOFs and quantify their chemical properties.
+S.C.
+provided expertise on metrics to be used to study the chemical properties and stability of MOFs with AI tools developed in this manuscript.
+D.C.
+provided expertise on expected features of MOFs to be used for carbon capture at industrial scale.
+I.F.
+guided the development of our ready-to-use AI framework in modern computing environments, and provided guidance on how to use and interpret metrics for MOF synthesizability.
+E.T.
+provided expertise on MD, chemistry and synthesis pathways of MOFs as well as how to analyze MD results.
+All authors reviewed and contributed to the manuscript.
+Corresponding author
+Correspondence to Eliu Huerta.
+Ethics declarations
+Competing interests
+The authors declare no competing interests.